### v2

# Training Language Models on Synthetic Edit Sequences Improves Code Synthesis

Abstract
Software engineers mainly write code by editing existing programs. In contrast, large language models (LLMs) autoregressively synthesize programs in a single pass. One explanation for this is the scarcity of open-sourced edit data. While high-quality instruction data for code synthesis is already scarce, high-quality edit data is even scarcer. To fill this gap, we develop a synthetic data generation algorithm called LintSeq. This algorithm refactors existing code into a sequence of code edits by using a linter to procedurally sample across the error-free insertions that can be used to sequentially write programs. It outputs edit sequences as text strings consisting of consecutive program diffs. To test LintSeq, we use it to refactor a dataset of instruction + program pairs into instruction + program-diff-sequence tuples. Then, we instruction finetune a series of smaller LLMs ranging from 2.6B to 14B parameters on both the re-factored and original versions of this dataset, comparing zero-shot performance on code synthesis benchmarks. We show that during repeated sampling, edit sequence finetuned models produce more diverse programs than baselines. This results in better inference-time scaling for benchmark coverage as a function of samples, i.e. the fraction of problems “pass@k” solved by any attempt given “k” tries. For example, on HumanEval pass@50, small LLMs finetuned on synthetic edit sequences are competitive with GPT-4 and outperform models finetuned on the baseline dataset by +20% ($\pm$3%) in absolute score. Finally, we also pretrain our own tiny LMs for code understanding. We show that finetuning tiny models on synthetic code edits results in state-of-the-art code synthesis for the on-device model class. Our 150M parameter edit sequence LM matches or outperforms code models with twice as many parameters, both with and without repeated sampling, including Codex and AlphaCode.
1 Introduction
The successes of large language models (LLMs) are difficult to overstate. However, consistent and correct zero-shot generation in code synthesis remains out-of-reach for all but the largest models (Abdin et al., 2024; Groeneveld et al., 2024; Dubey et al., 2024). Compared to other reasoning tasks, this setting has two challenging properties, namely solutions are both structured and long-form.
Humans tackle problems that have these properties by leveraging abstract mental models, first developing a plan for their solution that reflects the setting’s structure and then executing the plan one step at a time (Gopnik, 1982; Kirsh, 2009). For example, a software engineer might employ object-oriented programming when creating a new code-base by developing a “class” object and then gradually adding new functionality to this class as their code-base becomes more complex.
Figure 1: Code synthesis with LMs trained on synthetic code edit sequences. Left: An example generation from an LM trained to synthesize code as a stream of static-error-free edits. Right: Comparing zero-shot HumanEval coverage (in %) as a function of FLOPs for large external LLMs vs the smaller LMs that we instruction finetune in this paper. Repeatedly sampling from edit sequence LMs yields coding problem solutions that are competitive with GPT-4 and GPT-4-Omni, and have total cost similar to sampling once from the best open-source LLMs (see Appendix F.4).
In contrast, LLMs are trained to autoregressively synthesize entire programs from scratch. This makes repeatedly editing a program with an LLM extremely expensive – current state-of-the-art, LLM-powered code editing tools like Cursor repeatedly prompt models to rewrite entire programs during every edit generation call (Sanger, 2024). LLM outputs also suffer from degrading quality as sequence lengths grow and exhibit limited diversity across samples (Chen et al., 2021; Li et al., 2022b; Roziere et al., 2023; Lozhkov et al., 2024). The consequence of these pathologies is that there does not exist a reliable trade-off between zero-shot generation quality and inference-time compute cost under the current paradigm of autoregressive code synthesis, particularly for smaller models.
In this paper, we claim that these issues can be mitigated at the data-level by reparameterizing code synthesis as a sequential edit problem. Rather than training models for single-step generation of entire programs, we propose that models be trained to generate code by predicting code edit sequences.
This objective has a major obstacle: while high-quality solution data for code synthesis is scarce, open-source edit data with good coverage over the distribution of all error-free code “diffs” is non-existent (Muennighoff et al., 2023). To address this, we introduce an algorithm titled “LintSeq” that re-factors existing programs into sequences of static error-free code edits. LMs trained on data generated with our algorithm output code edits that effect interdependent lines of a program. LintSeq is parameter-free. It consists of two phases: a backward sampling phase, which takes a source file as input and uses a static program verifier to sample sequences of error-free program states that begin with this file and end with empty programs; and a forward edit computation phase, which reverses each sequence of programs, employing the Unix diff (Thompson & Ritchie, 1975) operator to compute deltas between consecutive versions of the source file, and outputs edit sequences. The static program verifier used by LintSeq in its backward sampling phase is also referred to as a linter.
To test the impact of finetuning LMs to synthesize code with edit sequences via LintSeq instruction data, we conduct a series of experiments comparing these models to those finetuned on a standard version of the same instruction data. We evaluate LMs zero-shot on the code synthesis benchmarks HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) by computing coverage, the proportion of problems solved by any attempt, as a function of samples. We do not allow LMs to use chain-of-thought reasoning (Wei et al., 2022) during evaluation. Our results are five-fold:
  1. 1.
    Across models ranging in scale from 150M to 14B parameters, instruction finetuning LMs on edit sequences vs full programs improves the quality and diversity of synthesized code.
  2. 2.
    The improved diversity of samples means that pass@k performance increases smoothly as a function of inference-time compute, allowing for a better trade-off between the two.
  3. 3.
    Tiny edit sequence LMs have state-of-the-art performance in their model class (Table 1).
  4. 4.
    For smaller LLMs, repeatedly sampling from edit sequence models results in HumanEval coverage that is competitive with GPT-4 models and has similar cumulative cost to sampling once per problem from open-source LLMs like Llama 3.1 405B (Figures 1 and 4).
  5. 5.
    Ablating the linter from edit sampling during data generation hurts the downstream quality of programs synthesized by edit sequence models (Figures 5 and 6).
2 LintSeq: Code Synthesis as a Sequential Edit Problem
Figure 2: Visualizing LintSeq, an algorithm for re-factoring programs into sequences of edits. This algorithm samples across all of the static error-free line insertions that can be used to write a program chunk-by-chunk. It uses the Unix-diff operator to express generated edit sequences as text strings.
LintSeq is an algorithm for synthetic data generation that samples across sequences of insertions that can be used to sequentially write a program, leveraging a linter to repeatedly check program states for static errors during generation. A linter is a type of standard code analysis tool that verifies the correctness of programs. The LintSeq algorithm is loosely inspired by recent work on discrete diffusion methods for text generation, where decoding is non-autoregressive (Li et al., 2022a).
The key hypothesis underlying LintSeq is as follows: by training LMs on code edit sequences with teacher-forced supervised learning, we can potentially achieve a better trade-off between generation quality and compute at inference-time while still benefiting from the training and sampling efficiency of the autoregressive modeling paradigm. In this section, we provide a formalism for the edit sequence re-parameterization of code synthesis and we formally introduce LintSeq.
2.1 Reparameterizing Code Datasets with Edits
We operate in the textual supervised learning setting in this paper, where we have access to a code dataset $\mathcal{D}$ of $N$ example programs $y$, each of which may be optionally paired with a corresponding natural language instruction $x$ that describes the program’s function, i.e. $\mathcal{D}=\{(x^{i},y^{i})\}_{i=1}^{N}$.
Let $\Delta(\cdot,\cdot)$ denote the Unix diff operator (Thompson & Ritchie, 1975), which computes a text difference between a pair of strings by performing a line-by-line matching and returns a string summarizing the detected differences. The diff operator is implemented by popular version control and software development systems to help programmers track edits or code “diffs” between versions of text files. A single edit computed with the diff operator may consist of multiple line deletions and/or line insertions.
Fix a program $y$ in the dataset $\mathcal{D}$. Consider a sequence $\bm{\sigma}_{y}$ of $j$ text strings corresponding to programs or program states that terminates at $y$.
$\bm{\sigma}_{y}=(y_{1},\dots,y_{j-1},y)$ (1)
We can equivalently re-express $\bm{\sigma}_{y}$ as an edit sequence $\bm{\delta}_{y}$ of length $j$ by first computing a diff between an empty program $\epsilon$ and the first program in the sequence, and then computing diffs between all pairs of consecutive programs, as shown below.
$\bm{\delta}_{y}=(\Delta(\varepsilon,y_{1}),\Delta(y_{1},y_{2}),\Delta(y_{2},y_{3}),\dots,\Delta(y_{j-1},y))$ (2)
If $\mathcal{D}^{\prime}$ is a dataset such that for every pair $(x,y)\in\mathcal{D}$, there exists a pair $(x,\bm{\delta}_{y})\in\mathcal{D}^{\prime}$, then we say that $\mathcal{D}^{\prime}$ is an edit sequence re-factoring of $\mathcal{D}$.
Figure 3: Empirics of processing code data with LintSeq. Left: Lines per example in a dataset of instruction finetuning data for Python synthesis before and after processing with LintSeq via the linter pylint (see Section 3.2). LintSeq processing adds lines of diff metadata to examples (see Appendix A). Right: The corresponding edit counts per synthetic code edit sequence. On a dataset of short programs (14 lines of code, on average), the mean LintSeq edit sequence contains four edits.
2.2 Generating Linter-Guided Synthetic Edit Sequences
Recall from above that a single program edit computed by the diff operator $\Delta(\cdot,\cdot)$ can consist of any number of deletions and insertions. LintSeq is an algorithm for computing edit sequence re-factorings $\mathcal{D^{\prime}}$ such that all data $(x,\bm{\delta}_{y})\in\mathcal{D^{\prime}}$ have a particular property: every edit in $\bm{\delta}_{y}$ consists of insertions only. There are two phases in LintSeq: a backward sampling phase that is used to compute program state sequences $\bm{\sigma}_{y}$, and a forward edit sequence computation phase that is used to re-express $\bm{\sigma}_{y}$ as edit sequences $\bm{\delta}_{y}$. An illustration of these phases is shown in Figure 2. Full examples of edit sequences generated with LintSeq are provided in Appendix E (Figures 9 and 10).
Phase I: Backward Sampling
In the backward sampling phase of LintSeq, for each of the $N$ pairs $(x,y)\in\mathcal{D}$, we generate $s$ sequences of intermediate program states $\bm{\sigma}_{y}$ that begin with the empty program and terminate at the original program $y$. These sequences are generated in reverse or backwards using a simple procedure that we dub linter-guided sampling. Starting with the program $y$, we sequentially generate each predecessor program in $\bm{\sigma}_{y}$ from its successor by following these steps: (1) delete a line from the current program by sampling uniformly at random; (2) run a linter or other verifier on the remaining code; (3) if the deletion induced new errors, remove all affected lines; and (4) repeat steps 2 and 3 until no errors are caught by the linter. We repeat these steps until all lines have been removed from the original program $y$, at which point $\bm{\sigma}_{y}$ has been generated.
Phase II: Forward Edit Computation
Once $s$ program state sequences $\bm{\sigma}_{y}$ have been generated for each $(x,y)\in\mathcal{D}$, we run the forward edit computation phase of our algorithm. In this phase, we apply Equation 2 from above to compute an edit sequence $\bm{\delta}_{y}$ for each $\bm{\sigma}_{y}$. Starting from the last program that was added to $\bm{\sigma}_{y}$, we use the diff operator to compute edits between each pair of consecutive programs in $\bm{\sigma}_{y}$ up to the original program $y$. Finally, we pair each edit sequence $\bm{\delta}_{y}$ with its instruction $x$ (if present) to yield an edit sequence re-factoring $\mathcal{D}^{\prime}$ of $\mathcal{D}$ with size $sN$.
2.3 Properties of LintSeq Data
Synthetic edit sequences generated by LintSeq have a few other important properties. Let $\bm{\delta}_{y}$ be an arbitrary $j$-length edit sequence in $\mathcal{D}^{\prime}$ generated with LintSeq, $\bm{\delta}_{y}=(\Delta(\varepsilon,y_{1}),\dots,\Delta(y_{j-1},y))$. First, we observe that there is a simple correspondence between $\bm{\delta}_{y}$ and the original program $y$ used to generate it: $y$ can be re-constructed by starting with an empty program, and successively applying each edit in $\bm{\delta}_{y}$ to this program one-by-one. In other words, the edit sequence $\bm{\delta}_{y}$ resolves to $y$. Furthermore, by construction, every prefix subsequence of $\bm{\delta}_{y}$ resolves to a sub-program of $y$ that is error-free, i.e. that throws no errors when checked with the linter or the verifier used during generation. These two properties, in conjunction with the uniform sampling step used in the first phase of the algorithm, show that LintSeq samples $s$ examples across all possible static error-free sequences of line insertions that can be used to sequentially write a program $y$ from-scratch.
We show an example of program synthesis dataset statistics before and after LintSeq processing in Figure 3. In the worst case, re-expressing a program as an edit sequence increases the length of a training example by a token count that is constant in the number of program lines (Appendix A).
2.4 Practicalities of Training Language Models on LintSeq Data
LintSeq can be run on any code data. It is agnostic to the contents of a program, and only depends on knowledge of the programming language that a file is written in and the existence of a linter or another kind of verifier for written program files.
We use teacher-forced supervised learning (Williams & Zipser, 1989) to train models on LintSeq data, concatenating edit sequences into a single string by interleaving edits with special tokens, “<|diff|>,” and computing instruction-conditioned losses over the resultant sequences. During inference, finetuned models can be prompted to synthesize programs with edit sequences by appending these special tokens to the ends of prompts. More details are provided in Appendix A.
Synthetic data generation with LintSeq is controlled by a single hyperparameter: the number of edit sequences $s$ that are sampled for each example in the source code dataset $\mathcal{D}$. Edit sequence sampling can optionally be constrained to avoid repetitions, though this may yield a dataset $\mathcal{D}^{\prime}$ with fewer than $s|\mathcal{D}|$ examples if there are programs in $\mathcal{D}$ with no non-empty and static error-free sub-programs.
3 Experiments
To study LintSeq and the impact of re-parameterizing program synthesis as a sequential edit generation problem, we conduct a multi-pronged set of instruction finetuning experiments. These experiments study code synthesis in Python and are designed to answer the following questions:
  • How does finetuning tiny LMs on re-factorized code edit data generated with LintSeq impact benchmark coverage compared to finetuning on the original code data?
  • Do performance improvements hold across model scales, families, and tokenizers?
  • How does ablating linter-guidance from LintSeq to finetune on randomly sampled edit sequences impact code synthesis?
Similar to previous works (Chen et al., 2021), we evaluate models by computing zero-shot coverage statistics on code synthesis benchmarks with and without repeated sampling. We expound upon our motivation for evaluating models in this manner in Appendix B.3.1.
3.1 Pretraining Tiny LMs for Code Understanding
We begin our investigations by pre-training two tiny decoder-only transformers, TinyCodeLM-150M and TinyCodeLM-400M, for Python code understanding on 72 billion tokens of text. Pretraining our own language models grants us a data contamination-free test-bed to study code synthesis with edit sequences, rapidly evaluate LintSeq, and broadly re-examine the trade-off between inference-time compute and generation quality in code synthesis for models that can be updated on-device.
We rely on open-source data and libraries to pretrain our models (Penedo et al., 2024; Lozhkov et al., 2024; Soldaini et al., 2024; Groeneveld et al., 2024). Our pretraining data mix is inspired by Code Llama (Roziere et al., 2023), and reflects a code-skewed mixture of web text and raw Python sampled from FineWeb and TheStack, respectively (Penedo et al., 2024; Li et al., 2023). The architecture of our models respectively mimics the two smallest versions of GPT-2 (Radford et al., 2019), but integrates the transformer architecture changes proposed by the OLMo framework. This includes the absence of bias terms and the addition of non-parametric layer norms (Ba, 2016), as well as the use of SwiGLU (Shazeer, 2020), rotary positional embeddings (Su et al., 2024), and the GPT-NeoX-20B tokenizer (Black et al., 2022). We train both models for two epochs with a batch size of 524,288 tokens on an NVIDIA H100 node with four GPUs. Our experiments are supported by Pytorch FSDP (Zhao et al., 2023). More details on our pretraining procedures are in Appendix C.
3.2 Generating a Synthetic Dataset with LintSeq
Next, to support our finetuning experiments, we prepare a large “baseline” dataset of paired instruction and program data. We re-factorize the programs in this dataset into code edit sequences with LintSeq.
Table 1: Summary of temperature-tuned coding benchmark results for LMs with $\mathbf{\leq 0.2}$B parameters. Scores annotated with “($\dagger$)” indicate external model evaluations that we ran using the procedure described in Appendix B, and all other scores are as reported by model authors. We list models in order of increasing HumanEval pass@1 score.
HumanEval MBPP(+)
Model Size pass@1 pass@10 pass@1 pass@10 Open-Source
PolyCoder 160M 2.1 3.4 - - 🌑
AlphaCode 89M 4.3 12.2 - - 🌕
Codex 85M 8.2 12.8 - - 🌕
SmolLM-Instruct 135M 7.6 14.4($\dagger$) 10.1($\dagger$) 14.6($\dagger$) 🌑
TinyCodeLM-Instruct 150M 9.1 13.2 11.5 16.4 🌑
TinyCodeLM-LintSeqInstruct 150M 12.8 20.6 13.6 24.4 🌑
Table 2: Summary of temperature-tuned coding benchmark results for $\mathbf{0.2}$B to $\mathbf{0.4}$B parameter language models. Annotations, model order, and evaluation procedure are the same as in Table 1.
HumanEval MBPP(+)
Model Size pass@1 pass@10 pass@1 pass@10 Open-Source
PolyCoder 400M 3.0 5.3 - - 🌑
TinyCodeLM-Instruct 400M 9.6 18.5 15.5 22.2 🌑
SmolLM-Instruct 360M 11.3 19.3($\dagger$) 19.4($\dagger$) 23.1($\dagger$) 🌑
AlphaCode 302M 11.6 18.8 - - 🌕
CodeT5+ 220M 12.0 20.7 - - 🌑
Codegen-Mono 350M 12.8 23.1 9.4($\dagger$) 15.2($\dagger$) 🌑
Codex 300M 13.2 20.4 - - 🌕
TinyCodeLM-LintSeqInstruct 400M 13.4 20.9 19.4 29.9 🌑
To that end, we first pool the Python portions of two open-source instruction datasets for code synthesis: the GPT 3.5/4-based Magicoder instruction tuning dataset and the StarCoder2-15B-based self-alignment training dataset (Wei et al., 2024b; a). These datasets are generated with the OSS-Instruct approach by Wei et al. (2024b) and have undergone decontamination for the benchmarks that we evaluate on in this paper. We conduct de-duplication on the pooled data to check for repeated examples. Furthermore, we strip any chain-of-thought-like natural language explanations from completion data. The resultant dataset has over 88,900 instruction-Python program pairs.
With our baseline dataset prepared, we run LintSeq to generate $s=5$ synthetic edit trajectory samples for each instruction-program pair. As described in Section 2.4, we concatenate each synthetic edit trajectory into a single string by interleaving consecutive edits with a special reserved “edit” token. Inspired by Muennighoff et al. (2024), we do not restrict against edit sequence repetitions. We use the popular Python linter pylint to guide edit sampling during generation. Examples of generated edit sequences and experiments testing the effect of varying $s$ are in Appendix E.
3.3 Finetuning Language Models on LintSeq Edit Sequences
We now probe the impact of instruction finetuning a variety of autoregressive LMs to synthesize code with edit sequences, compared to standard generation of full programs. We use two code synthesis benchmarks to support our model evaluations: HumanEval (Chen et al., 2021) and Mostly Basic Programming Problems (MBPP) (Austin et al., 2021).
Using both the code edit re-factorized and baseline instruction datasets obtained in section 3.2, we run pairs of finetuning experiments with six different models. In each experiment pair, we finetune an LM on both datasets for an equal number of optimizer steps and with the same learning rate schedule, saving intermediate checkpoints throughout finetuning. We run full benchmark evaluations on HumanEval and MBPP on each saved checkpoint [^1] To process the generations of edit sequence LMs into executable programs, we simply resolve each of the predicted code edits one-by-one. This procedure is visualized in Figure 1 and described in Appendix A.2., performing no prompt tuning. Then, we compare the best inference-time scaling behavior, i.e. benchmark coverage (pass@k) as a function of samples (k), obtained by edit sequence finetuning vs standard finetuning for each model. A more detailed description of the computed metrics as well as a full specification of the evaluation and finetuning procedures is provided in Appendices B and D.
Figure 4: HumanEval and MBPP coverage with repeated sampling (“pass@k vs. k”) achieved by instruction finetuning Gemma 2, Phi-3, and Llama 3.1 language models on a dataset of LintSeq edit sequence re-factored vs standard Python code (temperature $=1$, top-p $=0.95$).
3.3.1 TinyCodeLM
We run our first two pairs of finetuning experiments on TinyCodeLM-150M and TinyCodeLM-400M. These models were not pretrained on code synthesis instruction data, nor were they pretrained on any “diff”-like edit data. Our experimental results are summarized in Tables 1 and 2, where we compare the temperature-tuned performance of our models to the reported benchmark coverage of existing code LMs at similar parameter scales. We also report the inference-time scaling of benchmark coverage as a function of samples for our finetuned models in Appendix Tables 8 and 9.
For both the 150M and 400M parameter versions of TinyCodeLM, we find that finetuning LMs to synthesize code with edits via LintSeq data dramatically improves benchmark performance compared to baseline finetuning. Indeed, the edit sequence variants of TinyCodeLM outperform all existing code language models of comparable scale that we are aware of, including AlphaCode (Li et al., 2022b), Codex (Chen et al., 2021), CodeT5+ (Wang et al., 2023b), and the recent SmolLM-Instruct (Ben Allal et al., 2024). Our smaller edit sequence-finetuned model is particularly strong for its size, roughly matching or out-performing models with twice as many parameters including the 300M parameter version of Codex and the 302M-parameter version of AlphaCode (Tables 1 and 2).
3.3.2 Gemma 2, Phi-3, and Llama 3.1
The results above raise a natural question: do performance improvements from finetuning LMs to synthesize code with edit sequences hold for other model scales, architectures, and tokenizers? To test this, we conduct four additional pairs of instruction finetuning experiments on LMs from three model families, Gemma 2, Phi-3, and Llama 3.1, ranging in size from 2.6B to 14B. We employ pretrained-only model weights, if available. The results of these experiments are in Figure 4, where we plot zero-shot benchmark coverage as a function of samples for instruction finetuned models. Raw coverage scores are reported in Appendix Tables 10 and 11.
Most of our findings echo those of Section 3.3.1. Aggregating across sample counts, we find that finetuning models to synthesize code with edits improves overall zero-shot performance on HumanEval and MBPP compared to finetuning on the original data. This suggests that re-factoring code with edit sequences is an architecture- and tokenizer-independent mechanism for improving downstream LM outputs. Furthermore, as shown in Figure 4 and Tables 10 and 11, we find that the degree by which edit sequence LMs outperform baseline model variants increases with repeated sampling for all tested models, culminating in an average absolute gain in pass@50 of +20% ($\pm$ 3%) on HumanEval and +12% ($\pm$ 2%) on MBPP. This observation confirms the hypothesis posed in Section 2, showing that training LMs to synthesize code with edits using LintSeq data improves the relationship between cumulative inference-time compute and zero-shot performance.
At pass@1, however, our results are slightly more mixed than in Section 3.3.1. For Phi-3 models, we observe either no difference or a decrease in score between each pair of model-data variants. One explanation for this is bias: the Phi-3 models have been previously instruction finetuned and were likely to have been trained on standard code data, putting LM variants finetuned to generate edit sequence re-factorized code at a comparative disadvantage due to distributional shift.
Figure 5: HumanEval and MBPP “pass@k vs. k” achieved by finetuning TinyCodeLM models on linter-guided vs randomly sampled code edit sequences (temperature $=1$, top-p $=0.95$).
3.4 Ablating Linter-Guidance
The backward sampling phase of LintSeq uses a linter or other verifier to group interdependent lines of a program together during edit sequence generation, ensuring that all code edits resolve to programs that are free of static errors. We conclude our experiments by testing the importance of this design choice with TinyCodeLM models.
To do this, we replace the backwards procedure described in Section 2.2 with exclusively random sampling; during each step of the algorithm, we first sample the number of lines to delete from the current program uniformly at random, before sampling a set of lines with the desired count. Using this linter-ablated version of the algorithm, we generate a new synthetic edit sequence dataset with the same size as the LintSeq dataset used in all previous finetuning experiments, i.e. with $s=5$ example sequences per sample in the source dataset. The average number of edits per example in this dataset ($\overline{E}_{\text{RandSeqInstruct}}=3.9$) is empirically similar to its linter-guided counterpart ($\overline{E}_{\text{LintSeqInstruct}}=3.8$, see Figure 3). We employ the same procedure as the one used in Section 3.3 to instruction finetune TinyCodeLM models on the dataset of randomly sampled edit sequences.
In Figure 5, we compare the temperature $1$ inference-time scaling laws on HumanEval and MBPP obtained by finetuning models on randomly sampled vs static error-free edit sequences. Raw model scores are also provided in Appendix F, Tables 8 and 9. Ablating linter-guidance results in a decline in benchmark coverage. On HumanEval, linter ablation reduces absolute pass@50 score on TinyCodeLM-150M ($22.6\%\mapsto 17.7\%$) and on TinyCodeLM-400M ($26.8\%\mapsto 22.0\%$). MBPP pass@50 is similarly affected, dropping for both models ($34.5\%\to 30.2\%$ and $39.6\%\mapsto 34.5\%$). These results suggest that the error-free nature of edits in LintSeq instruction finetuning data does indeed have a positive impact on the coding problem coverage of sampled solutions.
To conclude our analysis, we probe whether training models on error-free edit sequences also has an effect on the presence of errors in across all generated programs, aside from its positive effects on coverage. To assess this, we run the Python linter pylint over each of the synthesized programs used to compute the reported temperature $1$ pass@k metrics, checking code for static errors.
In Figure 6, we plot the total proportions of synthesized program samples with at least one static error across finetuned model variants. On both benchmarks, LMs trained on randomly sampled edits (dark grey) appear to generate “buggy” code with much higher frequency than all other models. Furthermore, on HumanEval, we find that LintSeq models (indigo) synthesize programs with static errors at a higher frequency than baseline models (light grey), despite their higher coverage of benchmark coding problems. This additional finding suggests that model performance gains from LintSeq cannot simply be attributed to improvement in static error frequency across code – training on re-factored code must be helping models write generally better, more diverse programs.
In summary, the error-free nature of the linter-guided edits sampled in LintSeq appears to indeed be important for improving both the quality and diversity of sampled programs (Figure 5), as well as the overall correctness (Figure 6) of code synthesized by language models trained on edit sequences.
Figure 6: Comparing static error frequency in synthesized code samples across baseline vs edit sequence instruction finetuned model variants (n $=50$, temperature $=1$, top-p $=0.95$).
4 Related Work
Foundation Models for Code
Code synthesis is one of the oldest problems in computer science. Neural language model-based approaches such as Codex, AlphaCode, CodeT5+, CodeGen, StarCoder, and Code Llama have recently proven to be extremely competitive with previous methods (Chen et al., 2021; Li et al., 2022b; Wang et al., 2023b; Nijkamp et al., 2022; Li et al., 2023; Roziere et al., 2023). Today, foundation models trained on web text and code data dominate, and LLM-powered code editing tools like Github Copilot and Cursor are used by thousands of engineers every day (Heaven, 2024). Many general-purpose LLMs are also trained on code data. While the largest of these LLMs show strong performance on coding benchmarks, generations continue to suffer from limited meaningful output diversity, prompt sensitivity, and degrading quality on long-contexts (Achiam et al., 2023; Gemini Team et al., 2023; Dubey et al., 2024). Smaller models also lag behind (Abdin et al., 2024; Gemma Team et al., 2024; Ben Allal et al., 2024). As of the writing of this paper, directly prompting LLMs to generate code “diffs” results in low quality edits across models (Sanger, 2024). We claim that this is the result of a data problem and we attempt to address it in this work.
Finetuning on Synthetic Data
LLM post-training methods like supervised finetuning have been shown to be extremely powerful for improving model performance across tasks (Wei et al., 2021). However, high-quality datasets of paired instruction-response examples are extremely expensive to curate. One possible solution lies in synthetic data generation methods like Self-Instruct, wherein an LLM is prompted to generate instructions and/or responses from examples (Wang et al., 2022). Such data have been used extensively for improving LLM performance through self-refinement and/or knowledge distillation on coding tasks (Chaudhary, 2023; Roziere et al., 2023; Abdin et al., 2024; Lozhkov et al., 2024). We employ post-processed instruction data for code synthesis created with a method from this family, OSS-Instruct (Wei et al., 2024b), as the base of our experiments on re-factorizing code with code edit sequences via LintSeq. Unlike Self-Instruct-like synthetic data generation methods, our algorithm does not employ an LLM for data generation, and instead generates examples of error-free edit sequences from existing code data by using a simple linter.
Finetuning on Edits
Several works have investigated finetuning code LLMs on edit data. Notably, Muennighoff et al. (2023) instruction tune models on a 4TB dataset of GitHub commits pairing code changes with human instructions. Relatedly, Li et al. (2024) use GitHub commit data sourced from Python repositories to generate code editing instruction data with GPT 3.5/ChatGPT. Both of these works specifically focus on better-equipping LLMs for natural language-prompted code editing tasks, in which a model is explicitly prompted to generate an edit in response to a natural language specification. Our work differs in three important ways: first, we study edit sequences rather than single edits; second, we train LLMs to predict edits implicitly during code synthesis; third, our synthetic edit generation algorithm does not rely on the existence of any kind of commit data.
“On Device” Language Models
As the capabilities of LLMs have improved, so to have those of small language models. Recent projects like SmolLM (Ben Allal et al., 2024) and OpenELM (Mehta et al., 2024) re-examine the potential of tiny language models that can be run and even updated “on-device,” i.e. on a smart phone or laptop. The representations learned by such models during pretraining are weaker than those of scaled-up LLMs (Kaplan et al., 2020). This is particularly true for harder tasks that involve reasoning, such as code synthesis (Gemma Team et al., 2024; Abdin et al., 2024). To our knowledge, the most recent open-source work studying small language models pretrained entirely for code understanding is from several years ago (Xu et al., 2022; Nijkamp et al., 2022; Wang et al., 2021; 2023b). The 150M and 400M parameter TinyCodeLM models pretrained in this paper belong to the “on device” model family and build upon previous works. These models provide an efficient test-bed for experiments on LM code synthesis that is updated to recent advancements in high throughput pretraining and to improvements in open-source data quality.
Inference-Time Compute Scaling
The performance of language models can be boosted during inference by using scaled-up sample counts, hand-engineered prompting schema, and/or search (Brown et al., 2024; Snell et al., 2024). These methods dramatically increase inference costs. Their effectiveness is tightly linked to the expressivity of learned model representations and the diversity of outputs across samples. Our experiments with smaller language models are inspired by these works – we study whether it is possible to (1) improve the expressivity of representations for code synthesis across LM parameter scales during finetuning, and (2) take advantage of this property to improve the inference-time performance of smaller LMs by larger margins during repeated sampling.
5 Discussion, Limitations, and Conclusion
This paper introduces an algorithm, LintSeq, for generating synthetic code edit sequences from existing programs. LintSeq enables LLM reasoning settings like code synthesis to be re-parameterized at the data-level as sequential edit generation tasks. The algorithm is parameter-free, requires only CPU to run, and makes no assumptions about the content or structure of code files.
Re-parameterizing code generation with edits has a few immediate benefits. For example, it makes code generation with LLMs much more controllable at the prompt-level (Appendix A.3) and it reduces the cost of predicting useful and correct code insertions with models, since synthetic edit-trained LLMs do not need to be prompted to re-synthesize entire programs from scratch (Section 2.4).
In our experiments with LintSeq, we also show the following:
  1. 1.
    Tiny LMs can be efficiently finetuned to synthesize Python programs with edit sequences via LintSeq data. This results in state-of-the-art code benchmark performance for models that can be trained “on device” (Sections 3.1 and 3.3.1).
  2. 2.
    Across other tested models from the Phi, Gemma, and Llama families, finetuning on LintSeq data also improves the diversity of zero-shot generations, boosting the inference-time scaling of coverage on HumanEval and MBPP at fixed sample counts (Section 3.3.2).
  3. 3.
    On HumanEval, the cumulative inference cost of repeatedly sampling from small edit sequence LLMs is similar to sampling once from larger LLMs and yields coverage that is competitive with GPT-4, GPT-4-Omni, and Llama 3.1 405B (Figure 1, Appendix F.4).
  4. 4.
    Ablating linter-guidance from LintSeq hurts the quality, diversity, and correctness of code synthesized by instruction finetuned models (Section 3.4).
There are several limitations to our work.
First, as currently formulated, LintSeq can only be used to generate synthetic sequences of insertion edits. This is a consequence of the parameter-free nature of the algorithm – every edit in a LintSeq sequence reflects an existing line of code in the source file used to generate it. As a result, LintSeq edit data can only be used to train models for synthesis-like tasks, rather than for refinement.
Second, our experiments with LintSeq study code synthesis in Python only. We look forward to testing LintSeq with other programming languages, verifiers, and problems in future work. One other exciting setting where LintSeq might improve the tradeoff between model generation quality and inference-time compute is in mathematical reasoning and formal theorem proving.
Ethics Statement
This work explores data-driven mechanisms for improving the quality of language model-generated code. Our synthetic data generation method relies on open-source data and our experiments leverage open-source software and resources. It is important to acknowledge that all language models for code synthesis have the potential to be misused – whether intentionally or unintentionally – for generation of code with vulnerabilities and/or malicious behaviors. Any and all model generated code has the potential to be harmful and must not be executed without precautions.
Acknowledgements
This work was supported by grants from NSF award 2339096 and ONR awards N00014-21-1-2758 and N00014-22-1-2773. We are grateful to Shenglong Wang and NYU High Performance Computing for their support of this project. UP is funded by an NSF GRFP Award, and LP is funded by the Packard Fellowship. We would like to thank Nate Rahn, Mahi Shafiullah, and David Brandfonbrener for helpful comments and discussions.
Appendix A More on Edit Sequences and Diffs
A.1 Reading Unix Diffs
We provide a guide to reading Unix-style diffs below in Figure 7. The diff shown in this figure is computed using the Python library difflib, which is the implementation that we use to compactly represent edits in our synthetic data generation experiments. Note that the total extra tokens present in an insertion edit sequence representation of a program scales with the number of program lines $L$, and can be upper-bounded as $T_{\text{diff}}\leq L\cdot((\text{chars in ``decorator''})+(\text{extra chars per line in ``body''}))=16L$.
Figure 7: The anatomy of a Unix diff: A diagrammatic visualization of the different parts of a Unix-style diff, as computed by difflib. The body of a diff can consist of multiple line deletions, followed by multiple line insertions. The decorator portion of the diff respectively indicates the location and size of these deletions and insertions, if any. Like the diff shown above, the edits in synthetic edit sequences generated by LintSeq consist of line insertions only.
A.2 Resolving Edit Sequences
During inference, LMs that have been finetuned on LintSeq instruct data will synthesize code via edit sequences, outputting text strings that consist of a sequence of consecutive Python diffs interleaved with newline characters and “<|diff|>” tokens, similar to  Piterbarg et al. (2024). Each of these diffs will be structured as shown in Figure 7, if correctly formatted by the language model.
Resolving an edit sequence generated by a language model into an executable Python program is simple: starting with an empty program, we consecutively apply the line insertions and/or deletions in the body of each diff to the lines of the program specified in its decorator. We continue this process until all of the diffs in the generated edit sequence have been parsed and resolved.
Figure 1 shows a code edit sequence generation from a LintSeq instruction finetuned LM and the corresponding resolved, executable Python program.
A.3 Controllability of Code Synthesis with Edit Sequence LMs
The structure of Unix-style diffs affects the downstream controllability of code synthesis with models that have been trained on edit sequence re-parameterized programs. As shown in Figure 7, the first line of every diff is a decorator that describes the location and the numbers of lines changed by the edit. During inference, autoregressive language models that have been trained on Unix-style diffs with this format can be prompted to predict an edit in any desired target location within the program being synthesized by “intervening” on a model generation.
A.4 Future Work: Searching in Edit Space
If we apply the lens of reinforcement learning or search to this setting, we might say that re-parameterizing the code data used to train a language model re-parameterizes the model’s generative action space. It is possible that combining edit sequence LMs with more sophisticated decoding mechanisms, inference-time search, and/or interactive post-training may result in even larger improvements to the quality of generated code than those of the zero-shot code synthesis settings studied in this paper. We look forward to testing this hypothesis in future work.
Appendix B Evaluation
HumanEval (Chen et al., 2021) and Mostly-Basic Programming Problems (MBPP) (Austin et al., 2021) are two of the most studied benchmarks for evaluating code LMs (Liu et al., 2023). These benchmarks probe the code synthesis capabilities of models, and consist of pairs of natural language program descriptions and test-cases. We employ the extended MBPP test cases released as MBPP(+) by Liu et al. (2023) to add additional rigour to our testing procedure. All of the code LMs that we compare our models against evaluate HumanEval performance using the original set of benchmark test cases; for consistency, we employ these same test cases in our evaluations when comparing the performance of our models to the reported scores of external LMs.
During testing on both HumanEval and MBPP(+), LMs are prompted to generate outputs using the natural language descriptions of target programs. Their outputs are then evaluated on the paired test cases. A generation is considered “correct” if and only if it passes all of the test cases upon execution, subject to a fixed timeout setting. Previous works on code synthesis with language models report scores on HumanEval and MBPP(+) across samples. The most common of these metrics is known as pass@k (Chen et al., 2021; Austin et al., 2021; Li et al., 2022b; Wang et al., 2023b). This is the metric that we use to report and compare model performance throughout this paper. Our implementation of HumanEval and MBPP(+) evaluations mimics the Big Code Evaluation Harness by Ben Allal et al. (2022). We do not allow models to use chain-of-thought during generation.
B.1 Prompting
The primary goal of this paper is to introduce a method for re-factorizing code synthesis with LMs by finetuning them on synthetic instruction data. As a result, we evaluate all models using minimal prompt formats, performing no prompt tuning (see Figures 9 and 10). Examples of the prompt formats that we use during evaluation are shown in Figure 8.
Figure 8: Examples of formatted HumanEval and MBPP(+) prompts used in model evaluations.
We finetune all tested models on example outputs exclusively corresponding to Python code, and as a result, we do not use Markdown formatting to separate Python code from natural language in either our instruction data nor in our inference-time prompts.
To evaluate models on HumanEval, we use both the default “Python version” prompt format in the original benchmark dataset, where a natural language program description is provided to an LM within a docstring, as well as the equivalent, fully natural language prompt format from HumanEvalPack (Muennighoff et al., 2023). The latter format is similar to the structure of the instructions in our finetuning datasets.
To evaluate models on MBPP(+), we use the default prompts from the MBPP benchmark dataset, formatted with specification of the target function name and arguments both inside and outside of the natural language instruction, as shown in Figure 8.
During LM benchmark evaluations, we test models on each of the format variants described above and report scores on the better performing variant for each benchmark only.
B.2 Generation and Parsing
During generation, we continue decoding until an end-of-sequence token is output by an LM. We treat all LM outputs as either Python code or sequences of Python code edits, depending on whether an LM was finetuned on standard instruct or LintSeq instruct data. In the latter case, we post-process outputs by resolving the output edit sequences using the procedure described in Appendix A.2.
B.3 Evaluating Model Checkpoints
B.3.1 Philosophy
There is a well-known trade-off between the temperature used for sampling from autoregressive code LMs and the benchmark coverage achievable by models, i.e. the proportion of problems “pass@k” for which an LM is able to generate at least one output that passes all test cases given “k” tries. This trade-off was first described by Chen et al. (2021). Informally, increasing the sampling temperature increases the width of the distribution from which tokens are sampled, producing more diverse but noisier (and possibly lower quality) generations. For larger repeated sample counts, the pass@k score typically increases with sampling temperature up to some threshold, beyond which the negative effects of noise overpower the positive effects of diversity. The benchmark coverage achievable by an LM at any temperature and in the limit of samples, i.e. on pass@k for $k\uparrow\infty$, ultimately depends on both the power and expressivity of the code language model’s learned representation.
[…]
B.3.2 Computing Coverage (Pass@K)
Our goal is to probe whether re-parameterizing code synthesis with edit sequences can improve the expressivity of smaller LLM representations, boosting benchmark coverage as a function of samples-per-problem. Hence, we primarily compare finetuned models by evaluating them with the procedures described above on HumanEval and MBPP(+) at a high temperature and a large sample count, computing pass@k for k$\ \in\{1,5,10,20,50\}$ with $N=50$ samples [^2] These are the largest sample counts that are feasible to compute on our hardware given the scope of our experiments. at temperature 1, top-p 0.95. We compute pass@k statistics with the same procedure as Chen et al. (2021). The results of these evaluations are reported throughout the paper and shown in Figures 4, 5 and Tables 8, 9, 10, 11. In each of these Figures and Tables, we identify the most performant checkpoint from each model-data finetuning run by comparing pass@50 score at temperature 1 on HumanEval and MBPP(+) across checkpoints.
Many existing state-of-the-art code synthesis LMs only report temperature-tuned pass@k scores on HumanEval, including Codex, AlphaCode, and Codegen-Mono (Chen et al., 2021; Li et al., 2022b; Nijkamp et al., 2022). Thus, in Tables 1 and 2, we select the single best overall checkpoint of each TinyCodeLM model by averaging HumanEval and MBPP(+) pass@50 score. Then, we temperature-tune each checkpoint’s pass@1 and pass@10 scores when reporting results. On HumanEval, we test temperatures $\tau\in\{0.0,0.2,0.4,0.8,1.0\}$. On MBPP(+), we sweep over a smaller temperature range, $\tau\in\{0.0,0.1,1.0\}$. We perform the same temperature tuning procedure when reporting external model benchmark scores as well, i.e. the scores annotated with “()($\dagger$)( † )” in Tables 1 and 2. When running benchmark evaluations with these external code LMs, we stray from the prompt formatting, generation, and parsing procedures described in Appendices B.1 and B.2; instead, in the interest of a fair evaluation, we reproduce the conventions reported by model authors to report other scores.
Appendix C Pretraining
We rely on data and libraries open-sourced by the HuggingFace, FineWeb, StarCoder, Dolma, OLMo, and PyTorch FSDP projects to pretrain our models (Wolf et al., 2020; Penedo et al., 2024; Lozhkov et al., 2024; Soldaini et al., 2024; Groeneveld et al., 2024; Zhao et al., 2023).
C.1 Model Architectures and Pretraining Hyperparameters
Table 3: Architectural and pretraining hyperparameters of our “on device” 150M and 400M parameter TinyCodeLM models, pretrained on a mixture of Web text and code for Python understanding.
TinyCodeLM
Smallest, 150M Parameters Small, 400M Parameters
Transformer Architecture decoder-only decoder-only
Model Family OlmoForCausalLM OlmoForCausalLM
Tokenizer GPT-NeoX-20B-OLMo GPT-NeoX-20B-OLMo
Attention Bias False False
Attention Dropout 0.0 0.0
Hidden Activation SwiGLU SwiGLU
Hidden Size 768 1024
Intermediate Size 3072 4096
Number of Attention Heads 12 16
Number of Hidden Layers 12 24
Number of Key-Value Heads 12 16
Vocabulary Size 50304 50304
Positional Encodings Rotary (RoPE) Rotary (RoPE)
Mixed Precision BFLOAT16 BFLOAT16
Weight Tying True True
Flash Attention 2 True True
Optimizer AdamW AdamW
Learning Rate 0.0003 0.0003
Weight Decay 0.01 0.01
Betas $(0.9,0.95)$ $(0.9,0.95)$
Epsilon 1.0e-05 1.0e-05
Learning Rate Scheduler cosine (with warmup) cosine (with warmup)
Number of Warm-Up Steps 100 100
Alpha-f ($\alpha_{f}$) 0.1 0.1
Total Epochs of Pretraining 2 2
C.2 Pretraining Data Mix
Table 4: Pretraining data mix used to train both TinyCodeLM models. Datasets were tokenized and prepared using HuggingFace and Dolma tooling (Wolf et al., 2020; Soldaini et al., 2024).
Pretraining Data Source Subset Tokens Documents
FineWeb (Penedo et al., 2024) 10BT Sample 10.4BT 14.9M
The Stack (Kocetkov et al., 2022) Python Only 61.8BT 24.2M
Appendix D Instruction Finetuning
D.1 Baseline Instruction Dataset
Table 5 displays the data sources that are used to prepare the dataset described in Section 3.2. These data are pooled and preprocessed into instruction-program pairs by stripping away Markdown formatting and natural language explanations from completions (Figure 9 and 10). In our experiments, we use the resultant data to finetune baseline models, comparing their performance to those of LMs finetuned on edit sequences generated with LintSeq from the same set of instruction-program pairs.
HuggingFace Instruction Data Source Subset Examples
bigcode/self-oss-instruct-sc2-exec-filter-50k Full 50,661
ise-uiuc/Magicoder-OSS-Instruct-75K Python 38,284
88,945
Table 5: Instruction data mix used to prepare the baseline instruction dataset in Section 3.2.
D.2 Procedures and Hyperparameters
We instruction finetune all models with Microsoft DeepSpeed using the ZeRO++ protocol for stage three sharding. For the largest of these models, we also use CPU parameter offloading to accelerate experiments (Wang et al., 2023a; Ren et al., 2021). When finetuning models on LintSeq data, we add a new token “<|diff|>” to tokenizers (Section 2.4) and resize model embeddings accordingly.
In our experiments with Gemma 2, Phi-3, and Llama 3.1 models, we use HuggingFace to access and load pretrained model weights and tokenizers. As mentioned in the main body of the paper, we instruction finetune pretrained-only weights if open-sourced and available. This is the case for Gemma 2 and Llama 3.1 only, as of the writing of this paper.
Across all of the finetuning experiments conducted in this paper, we train model-data variants with the same batch size and for an equal number of total optimizer steps. This optimizer step count corresponds to ten epochs of finetuning with the baseline instruction tuning dataset described in Section 3.2. We save intermediate checkpoints at equal optimizer step intervals in all experiments.
[…]
TinyCodeLM Gemma 2 Phi-3 Llama 3.1
150M 400M 2B 3.8B 14B 8B
Peak Learning Rate ($\alpha$) 3e-4 3e-4 5e-5 5e-5 1e-5 1e-5
Table 6: Peak learning rates used to instruction finetune models.
Hyperparameter Setting
Learning Rate Scheduler linear
Warmup Ratio 0.001
Weight Decay 0.01
Total Batch Size 512
Batch Loss Reduction sum
Mixed Precision BFLOAT16
Max Sequence Length 1024
Total Optimizer Steps 1740
Table 7: All other instruction finetuning settings, re-used across experiments.
Appendix E More on Synthetic Data Generation with LintSeq
E.1 Examples of Generated Synthetic Edit Trajectories
Figure 9: LintSeq edit sequence samples vs baseline instruction-program data, example A.
Figure 10: LintSeq edit sequence samples vs baseline instruction-program data, example B.
E.2 Tuning LintSeq Example Count
Figure 11: Probing the effect of varying the number of edit sequences sampled with LintSeq per instruction-example pair during data generation: Using the source dataset described in Section 3.2, we sweep over the value of the LintSeq parameter $s$ used during synthetic data generation to yield three different edit sequence instruction datasets with $s\in\{1,5,10\}$. We finetune TinyCodeLM models on each of these datasets, and compare the resultant HumanEval and MBPP(+) performance vs samples (i.e. pass@k vs k) at temperature 1. The most performant values is $s=5$.
Appendix F Additional Results
F.1 Pretraining TinyCodeLM
Figure 12: Evaluating the zero-shot Python synthesis capabilities of TinyCodeLM-150M during pretraining on HumanEval and MBPP(+).
Figure 13: Evaluating the zero-shot Python synthesis capabilities of TinyCodeLM-400M during pretraining on HumanEval and MBPP(+).
F.2 Finetuning TinyCodeLM
F.2.1 Inference-Time Scaling Laws
Table 8: HumanEval fixed-temperature coverage scaling results achieved by finetuning TinyCodeLM models (zero-shot, temperature = 1, top-p = 0.95).
HumanEval
Model Variant Size Linter Guided pass@1 pass@5 pass@10 pass@20 pass@50
tinycodeLM-Instruct 150M - 6.2 10.4 12.3 14.5 18.3
tinycodeLM-RandSeqInstruct 150M 4.0 9.7 11.9 14.4 17.7 (-0.6)
tinycodeLM-LintSeqInstruct 150M 6.4 14.3 17.7 20.4 22.6 (+4.3)
tinycodeLM-Instruct 400M - 6.8 11.4 13.7 16.2 18.9
tinycodeLM-RandSeqInstruct 400M 7.2 12.7 15.6 18.6 22.0 (+3.1)
tinycodeLM-LintSeqInstruct 400M 7.2 14.9 18.2 21.5 26.8 ( +7.9)
Table 9: MBPP(+) fixed-temperature coverage scaling results achieved by finetuning TinyCodeLM models (zero-shot, temperature = 1, top-p = 0.95).
MBPP(+)
Model Variant Size Linter Guided pass@1 pass@5 pass@10 pass@20 pass@50
tinycodeLM-Instruct 150M - 7.3 16.4 20.4 24.6 30.6
tinycodeLM-RandSeqInstruct 150M 4.3 13.1 18.4 23.7 30.2 (-0.4)
tinycodeLM-LintSeqInstruct 150M 7.5 18.2 23.4 28.5 34.5 (+3.9)
tinycodeLM-Instruct 400M - 9.3 17.9 22.2 26.8 32.4
tinycodeLM-RandSeqInstruct 400M 5.5 15.9 21.5 27.0 34.5 (+2.1)
tinycodeLM-LintSeqInstruct 400M 13.1 25.2 29.9 34.4 39.6 (+7.2)
F.3 Finetuning Gemma 2, Phi-3, and Llama 3.1
F.3.1 Inference-Time Scaling Laws
Table 10: HumanEval fixed-temperature coverage scaling results achieved by finetuning Gemma 2, Phi-3, and Llama 3.1 models (zero-shot, temperature = 1, top-p = 0.95).
HumanEval
Model Variant Parameters pass@1 pass@5 pass@10 pass@20 pass@50
Gemma-2-Instruct 2.6B 16.1 26.3 30.5 34.6 40.2
Gemma-2-LintSeqInstruct 2.6B 21.6 35.0 41.2 48.3 58.5 (+18.3)
Phi-3-Mini-Instruct 3.8B 40.1 48.8 52.1 55.0 57.3
Phi-3-Mini-LintSeqInstruct 3.8B 38.3 63.6 72.0 79.0 86.6 (+29.3)
Llama-3.1-Instruct 8B 34.2 50.3 55.6 59.7 63.4
Llama-3.1-LintSeqInstruct 8B 37.6 61.1 69.2 75.3 82.3 (+18.9)
Phi-3-Med-Instruct 14B 52.8 67.9 72.0 75.0 77.4
Phi-3-Med-LintSeqInstruct 14B 44.7 73.5 80.2 84.5 90.2 (+12.8)
Table 11: MBPP(+) fixed-temperature coverage scaling results achieved by finetuning Gemma 2, Phi-3, and Llama 3.1 models (zero-shot, temperature = 1, top-p = 0.95).
MBPP(+)
Model Variant Parameters pass@1 pass@5 pass@10 pass@20 pass@50
Gemma-2-Instruct 2.6B 22.8 34.2 37.8 41.2 44.6
Gemma-2-LintSeqInstruct 2.6B 27.7 41.3 46.3 50.5 54.7 (+10.1)
Phi-3-Mini-Instruct 3.8B 35.2 45.9 49.1 52.0 56.1
Phi-3-Mini-LintSeqInstruct 3.8B 39.3 57.5 62.4 65.9 69.4 (+13.3)
Llama-3.1-Instruct 8B 38.4 50.2 53.8 56.6 59.4
Llama-3.1-LintSeqInstruct 8B 38.5 56.5 61.6 65.7 69.8 (+10.4)
Phi-3-Med-Instruct 14B 40.3 51.5 54.2 56.3 59.0
Phi-3-Med-LintSeqInstruct 14B 40.0 60.9 67.0 71.5 75.9 (+16.9)
F.4 Computing HumanEval Coverage vs Cumulative Inference-Time FLOPs
In Figure 1, we plot HumanEval coverage as a function of cumulative inference-time FLOPs, comparing the performance and total cost of repeatedly sampling from our instruction finetuned Phi-3 and Llama 3.1 models vs sampling a single generation per problem from very large models like Llama 3.1 405B (Dubey et al., 2024) and Nemotron 4 340B (Adler et al., 2024).
We use the approximations below, drawn from Kaplan et al. (2020), to conservatively estimate the cumulative inference costs of synthesizing solutions to all of the $164$ HumanEval benchmark problems across different models. The models that we compare are all dense transformers, where the majority of the parameters are used in matrix multiplications.
FLOPs per token $\approx 2\cdot(N_{\text{model-params}}+2\cdot L_{\text{model-layers}}\cdot C_{\text{context}})$
Total FLOPs $\approx\text{FLOPs per token}\cdot T_{\text{avg-total-tokens-per-sample}}\cdot K_{\text{samples}}\cdot M_{\text{problems}}$
$\approx\text{FLOPs per token}\cdot T_{\text{avg-total-tokens-per-sample}}\cdot K_{\text{samples}}\cdot 164$
For our instruction finetuned models, we determine the quantities $T_{\text{avg-total-tokens-per-sample}}$ by computing token counts over all sets of samples per problem that we obtained to compute the coverage statistics in Figure 4 and Table 10 above. These token statistics are provided in the table below.
Table 12: HumanEval total tokens per sample for finetuned Gemma 2, Phi-3, and Llama 3.1 models (zero-shot, temperature $=1$, top-p $=0.95$). These counts reflect prompt and completion tokens. They are computed using the same samples whose coverage statistics are reported in Table 10.
Model Variant Parameters Avg. Total Tokens Per HumanEval Sample ($n=50$)
Gemma-2-Instruct 2.6B 172
Gemma-2-LintSeqInstruct 2.6B 205
Phi-3-Mini-Instruct 3.8B 218
Phi-3-Mini-LintSeqInstruct 3.8B 247
Llama-3.1-Instruct 8B 178
Llama-3.1-LintSeqInstruct 8B 201
Phi-3-Med-Instruct 14B 230
Phi-3-Med-LintSeqInstruct 14B 247
Note that edit sequence (i.e. LintSeqInstruct finetuned) LMs have slightly higher average token counts per sample due to presence of “diff” descriptor tokens in generations (see Appendix A).
We report zero-shot HumanEval coverage for external models by using the evaluation statistics from Dubey et al. (2024) (Table 18, column one). To estimate cumulative inference-time FLOPs for these models, we employ the approximation expression above and estimate $T_{\text{avg-total-tokens-per-sample}}\approx 200$, reflecting an ensemble over the per sample token counts of standard “Instruct” finetuned models shown in Table 12. Note that this ensembled statistic reflects program generations without chain-of-thought only. As a result, we believe it to be a conservative estimator.

### v3

@@
 Software engineers mainly write code by editing existing programs.
-In contrast, large language models (LLMs) autoregressively synthesize programs in a single pass.
-One explanation for this is the scarcity of open-sourced edit data.
-While high-quality instruction data for code synthesis is already scarce, high-quality edit data is even scarcer.
+In contrast, language models (LMs) autoregressively synthesize programs in a single pass.
+One explanation for this is the scarcity of sequential edit data.
+While high-quality instruction data for code synthesis is scarce, edit data for synthesis is even scarcer.
 To fill this gap, we develop a synthetic data generation algorithm called LintSeq.
-This algorithm refactors existing code into a sequence of code edits by using a linter to procedurally sample across the error-free insertions that can be used to sequentially write programs.
-It outputs edit sequences as text strings consisting of consecutive program diffs.
-To test LintSeq, we use it to refactor a dataset of instruction + program pairs into instruction + program-diff-sequence tuples.
-Then, we instruction finetune a series of smaller LLMs ranging from 2.6B to 14B parameters on both the re-factored and original versions of this dataset, comparing zero-shot performance on code synthesis benchmarks.
-We show that during repeated sampling, edit sequence finetuned models produce more diverse programs than baselines.
-This results in better inference-time scaling for benchmark coverage as a function of samples, i.e.
-the fraction of problems “pass@k” solved by any attempt given “k” tries.
-For example, on HumanEval pass@50, small LLMs finetuned on synthetic edit sequences are competitive with GPT-4 and outperform models finetuned on the baseline dataset by +20% ($\pm$3%) in absolute score.
+This algorithm refactors programs into sequences of synthetic edits by using a linter to procedurally sample across interdependent lines of source code.
+Synthetic edits sampled with LintSeq reflect the syntax and semantics of their programming language.
+To test the algorithm, we use it to refactor a dataset of instruction + program pairs into instruction + program-diff-sequence tuples.
+Then, we fine-tune a series of smaller LMs ranging from 2.6B to 14B parameters on both the re-factored and original versions of this dataset.
+We perform comprehensive evaluations comparing edit sequence code LMs against baselines on HumanEval, MBPP(+), CodeContests, DS-1000, and BigCodeBench.
+We show that models fine-tuned to iteratively synthesize code match or outperform baselines on pass@1, and exhibit better scaling across higher pass@k as a function of total test-time FLOPs.
 Finally, we also pretrain our own tiny LMs for code understanding.
-We show that finetuning tiny models on synthetic code edits results in state-of-the-art code synthesis for the on-device model class.
-Our 150M parameter edit sequence LM matches or outperforms code models with twice as many parameters, both with and without repeated sampling, including Codex and AlphaCode.
+We show that fine-tuning these models to synthesize code edit-by-edit results in strong performance on HumanEval and MBPP(+) compared to existing code language models of similar scale such as CodeT5+, AlphaCode, and Codex.
 1 Introduction
-The successes of large language models (LLMs) are difficult to overstate.
+The successes of language models (LMs) are difficult to overstate.
 However, consistent and correct zero-shot generation in code synthesis remains out-of-reach for all but the largest models (Abdin et al., 2024; Groeneveld et al., 2024; Dubey et al., 2024).
-Compared to other reasoning tasks, this setting has two challenging properties, namely solutions are both structured and long-form.
+Compared to other reasoning tasks, this setting has two challenging properties, namely solutions are both long and structured.
 Humans tackle problems that have these properties by leveraging abstract mental models, first developing a plan for their solution that reflects the setting’s structure and then executing the plan one step at a time (Gopnik, 1982; Kirsh, 2009).
@@
 Figure 1: Code synthesis with LMs trained on synthetic code edit sequences.
-Left: An example generation from an LM trained to synthesize code as a stream of static-error-free edits.
-Right: Comparing zero-shot HumanEval coverage (in %) as a function of FLOPs for large external LLMs vs the smaller LMs that we instruction finetune in this paper.
-Repeatedly sampling from edit sequence LMs yields coding problem solutions that are competitive with GPT-4 and GPT-4-Omni, and have total cost similar to sampling once from the best open-source LLMs (see Appendix F.4).
-In contrast, LLMs are trained to autoregressively synthesize entire programs from scratch.
-This makes repeatedly editing a program with an LLM extremely expensive – current state-of-the-art, LLM-powered code editing tools like Cursor repeatedly prompt models to rewrite entire programs during every edit generation call (Sanger, 2024).
-LLM outputs also suffer from degrading quality as sequence lengths grow and exhibit limited diversity across samples (Chen et al., 2021; Li et al., 2022b; Roziere et al., 2023; Lozhkov et al., 2024).
-The consequence of these pathologies is that there does not exist a reliable trade-off between zero-shot generation quality and inference-time compute cost under the current paradigm of autoregressive code synthesis, particularly for smaller models.
+Left: An example generation from an LM trained to synthesize code as a stream of linter-error-free edits.
+Right: Training LMs to write code edit-by-edit by preprocessing instruction data for SFT with LintSeq improves test-time scaling laws during repeated sampling, i.e.
+the percentage of benchmark problems solved by any attempt (pass@k) as a function of total test-time FLOPs compared to training on standard data (see Appendix A.4).
+Shading indicates standard error in linear fit.
+In contrast, LMs are trained to autoregressively synthesize entire programs from scratch.
+This makes repeatedly editing a program with an LM extremely expensive – current state-of-the-art, LM-powered code editing tools like Cursor repeatedly prompt models to rewrite entire programs during every edit generation call (Sanger, 2024).
+LM outputs also suffer from degrading quality as sequence lengths grow and exhibit limited diversity across samples (Chen et al., 2021; Li et al., 2022b; Roziere et al., 2023; Lozhkov et al., 2024).
+The consequence of these pathologies is that there does not exist a reliable trade-off between zero-shot generation quality and total test-time compute under the current paradigm of autoregressive code synthesis, particularly for smaller LMs.
 In this paper, we claim that these issues can be mitigated at the data-level by reparameterizing code synthesis as a sequential edit problem.
-Rather than training models for single-step generation of entire programs, we propose that models be trained to generate code by predicting code edit sequences.
-This objective has a major obstacle: while high-quality solution data for code synthesis is scarce, open-source edit data with good coverage over the distribution of all error-free code “diffs” is non-existent (Muennighoff et al., 2023).
-To address this, we introduce an algorithm titled “LintSeq” that re-factors existing programs into sequences of static error-free code edits.
-LMs trained on data generated with our algorithm output code edits that effect interdependent lines of a program.
-LintSeq is parameter-free.
-It consists of two phases: a backward sampling phase, which takes a source file as input and uses a static program verifier to sample sequences of error-free program states that begin with this file and end with empty programs; and a forward edit computation phase, which reverses each sequence of programs, employing the Unix diff (Thompson & Ritchie, 1975) operator to compute deltas between consecutive versions of the source file, and outputs edit sequences.
-The static program verifier used by LintSeq in its backward sampling phase is also referred to as a linter.
-To test the impact of finetuning LMs to synthesize code with edit sequences via LintSeq instruction data, we conduct a series of experiments comparing these models to those finetuned on a standard version of the same instruction data.
-We evaluate LMs zero-shot on the code synthesis benchmarks HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) by computing coverage, the proportion of problems solved by any attempt, as a function of samples.
-We do not allow LMs to use chain-of-thought reasoning (Wei et al., 2022) during evaluation.
-Our results are five-fold:
+Rather than training models for single-step generation of entire programs, we propose that models be trained to generate code edit-by-edit.
+This objective has a major obstacle: while datasets of filtered GitHub repository commits like CommitPackFT (Muennighoff et al., 2023) have dramatically improved the quality of open-source code edit data, they contain limited sequential data.
+Moreover, the edits in such these datasets reflect the granularity at which programmers save code, but not necessarily the granularity at which they write and/or reason about it.
+To address this, we introduce a sampling algorithm called “LintSeq” that can be used to express any program in a training corpus as a sequence of structured code edits.
+LintSeq leverages linters – simple code analysis tools that check programs for errors and stylistic issues – to ensure that each generated edit meaningfully reflects the syntactical structure of the programming language that it is written in.
+The algorithm consists of two phases: a backward phase, which takes a source file as input and samples code deletions from this file to yield possible sequences of linter-error-free intermediate program states; and a forward edit computation phase, which reverses each sampled program sequence, employs the Unix diff (Thompson & Ritchie, 1975) operator to compute deltas between consecutive versions of each file, and outputs the generated edit sequences.
+LMs trained on data sampled with LintSeq synthesize code by repeatedly predicting insertion edits to files.
+To test the impact of training LMs on synthetic edit sequences sampled with LintSeq, we conduct a series of supervised fine-tuning (SFT) experiments.
+In each experiment, we compare the performance of models trained on a corpus of example programs re-sampled into synthetic edit sequences with LintSeq to those trained on the original dataset.
+We evaluate LMs zero-shot and without chain-of-thought on HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), DS-1000 (Lai et al., 2023), BigCodeBench (Zhuo et al., 2024), and CodeContests (Li et al., 2022b) on “pass@k,” the proportion of problems solved by any attempt given “k” tries.
+Our results show the following:
   1. 1.
-    Across models ranging in scale from 150M to 14B parameters, instruction finetuning LMs on edit sequences vs full programs improves the quality and diversity of synthesized code.
+    Across models ranging in scale from 150M to 14B parameters, training LMs to iteratively synthesize programs improves the diversity of model-generated code compared to training on standard instruction data, while either preserving or improving code quality.
   2. 2.
-    The improved diversity of samples means that pass@k performance increases smoothly as a function of inference-time compute, allowing for a better trade-off between the two.
+    The improved diversity of generated programs means that pass@k performance increases faster as a function of test-time FLOPs, allowing for a better trade-off between the two.
   3. 3.
-    Tiny edit sequence LMs have state-of-the-art performance in their model class (Table 1).
-  4. 4.
-    For smaller LLMs, repeatedly sampling from edit sequence models results in HumanEval coverage that is competitive with GPT-4 models and has similar cumulative cost to sampling once per problem from open-source LLMs like Llama 3.1 405B (Figures 1 and 4).
-  5. 5.
-    Ablating the linter from edit sampling during data generation hurts the downstream quality of programs synthesized by edit sequence models (Figures 5 and 6).
+    Ablating the linter from edit sampling during data generation hurts the downstream quality of programs synthesized by edit sequence models.
 2 LintSeq: Code Synthesis as a Sequential Edit Problem
-Figure 2: Visualizing LintSeq, an algorithm for re-factoring programs into sequences of edits.
-This algorithm samples across all of the static error-free line insertions that can be used to write a program chunk-by-chunk.
-It uses the Unix-diff operator to express generated edit sequences as text strings.
-LintSeq is an algorithm for synthetic data generation that samples across sequences of insertions that can be used to sequentially write a program, leveraging a linter to repeatedly check program states for static errors during generation.
-A linter is a type of standard code analysis tool that verifies the correctness of programs.
-The LintSeq algorithm is loosely inspired by recent work on discrete diffusion methods for text generation, where decoding is non-autoregressive (Li et al., 2022a).
-The key hypothesis underlying LintSeq is as follows: by training LMs on code edit sequences with teacher-forced supervised learning, we can potentially achieve a better trade-off between generation quality and compute at inference-time while still benefiting from the training and sampling efficiency of the autoregressive modeling paradigm.
-In this section, we provide a formalism for the edit sequence re-parameterization of code synthesis and we formally introduce LintSeq.
-2.1 Reparameterizing Code Datasets with Edits
+Figure 2: LintSeq: Training LMs to write code edit-by-edit with supervised learning by generating synthetic data.
+LintSeq decomposes existing programs into synthetic edits that reflect the syntax & semantics of their programming language.
+At each iteration, the algorithm samples an edit chunk from a program by: randomly selecting a line of code to delete; identifying the minimal set of lines that are dependent on this line with a code linter; and finally, removing the line and its dependents.
+These steps are repeated until all lines of code have been removed.
+LintSeq then processes the reversed sequence of program states with Unix-diff to express it as a sequence of edits.
+The key to solving a hard problem often lies in knowing how to decompose it into sub-problems.
+LintSeq is an algorithm for synthetic data generation that decomposes programs in training corpuses across insertion edit chunks that reflect the syntax and semantics of their programming language.
+To sample such chunks, it uses a code linter.
+The algorithm is inspired by recent work on discrete diffusion methods for text generation, where decoding is non-autoregressive (Li et al., 2022a).
+Informally, the hypothesis underlying LintSeq is as follows: by training LMs to synthesize code edit-by-edit on large-scale datasets, we can potentially achieve a better trade-off between generation quality and test-time compute while still benefiting from the training and sampling efficiency of autoregressive language modeling.
+In this section, we define important terms, provide a formalism for the edit sequence re-parameterization of code synthesis, and formally introduce LintSeq.
+2.1 Definitions
+We define a linter to be a static code analysis tool that scans source code for defects.
+Linters can identify code that is objectively incorrect, throwing errors if and when a program contains syntax errors or refers to non-existent variables or packages.
+It is important to note that unlike a formal verifier, linters may return false positives, i.e.
+they may be unable to detect more complex errors, particularly in dynamically typed programming languages like Python or JavaScript.
+For a given source file, define an intermediate program state to be a program that contains only a subset of the line-by-line contents of the original file, such that the order of these lines is preserved.
+We call an intermediate program state linter-error-free if checking this program with an appropriate linter produces exactly the same error trace(s) as those output when checking the original source file.
+2.2 Representing Code with Edit Sequences
 We operate in the textual supervised learning setting in this paper, where we have access to a code dataset $\mathcal{D}$ of $N$ example programs $y$, each of which may be optionally paired with a corresponding natural language instruction $x$ that describes the program’s function, i.e.
 $\mathcal{D}=\{(x^{i},y^{i})\}_{i=1}^{N}$.
-Let $\Delta(\cdot,\cdot)$ denote the Unix diff operator (Thompson & Ritchie, 1975), which computes a text difference between a pair of strings by performing a line-by-line matching and returns a string summarizing the detected differences.
-The diff operator is implemented by popular version control and software development systems to help programmers track edits or code “diffs” between versions of text files.
+Let $\Delta(\cdot,\cdot)$ denote the Unix diff operator (Thompson & Ritchie, 1975), which computes a text difference between a pair of strings by performing a line-by-line matching and returns a summary of the detected differences.
+The diff operator is implemented by popular version control and development systems to help programmers track edits between versions of text files.
 A single edit computed with the diff operator may consist of multiple line deletions and/or line insertions.
-Fix a program $y$ in the dataset $\mathcal{D}$. Consider a sequence $\bm{\sigma}_{y}$ of $j$ text strings corresponding to programs or program states that terminates at $y$.
-$\bm{\sigma}_{y}=(y_{1},\dots,y_{j-1},y)$ (1)
-We can equivalently re-express $\bm{\sigma}_{y}$ as an edit sequence $\bm{\delta}_{y}$ of length $j$ by first computing a diff between an empty program $\epsilon$ and the first program in the sequence, and then computing diffs between all pairs of consecutive programs, as shown below.
+Fix a program $y$ in the dataset $\mathcal{D}$.
+Consider a sequence of $\bm{\sigma}_{y}$ of $j$ text strings corresponding to programs that terminates at $y$, $\bm{\sigma}_{y}=(y_{1},\dots,y_{j-1},y)$.
+We can equivalently re-express $\bm{\sigma}_{y}$ as an edit sequence $\bm{\delta}_{y}$ of length $j$ by first computing a diff between an empty program $\varepsilon$ and the first program in the sequence, and then computing diffs between all pairs of consecutive programs, as shown below.
 $\bm{\delta}_{y}=(\Delta(\varepsilon,y_{1}),\Delta(y_{1},y_{2}),\Delta(y_{2},y_{3}),\dots,\Delta(y_{j-1},y))$ (1)
-If $\mathcal{D}^{\prime}$ is a dataset such that for every pair $(x,y)\in\mathcal{D}$, there exists a pair $(x,\bm{\delta}_{y})\in\mathcal{D}^{\prime}$, then we say that $\mathcal{D}^{\prime}$ is an edit sequence re-factoring of $\mathcal{D}$.
-Figure 3: Empirics of processing code data with LintSeq.
-Left: Lines per example in a dataset of instruction finetuning data for Python synthesis before and after processing with LintSeq via the linter pylint (see Section 3.2).
-LintSeq processing adds lines of diff metadata to examples (see Appendix A).
-Right: The corresponding edit counts per synthetic code edit sequence.
-On a dataset of short programs (14 lines of code, on average), the mean LintSeq edit sequence contains four edits.
+If $\mathcal{D}^{\prime}$ is a dataset such that for every pair $(x,y)\in\mathcal{D}$, there exists a pair $(x,\bm{\delta}_{y})\in\mathcal{D}^{\prime}$, then we say that $\mathcal{D}^{\prime}$ is an edit sequence refactoring of $\mathcal{D}$.
 2.3 Generating Linter-Guided Synthetic Edit Sequences
 Recall from above that a single program edit computed by the diff operator $\Delta(\cdot,\cdot)$ can consist of any number of deletions and insertions.
-LintSeq is an algorithm for computing edit sequence re-factorings $\mathcal{D^{\prime}}$ such that all data $(x,\bm{\delta}_{y})\in\mathcal{D^{\prime}}$ have a particular property: every edit in $\bm{\delta}_{y}$ consists of insertions only.
+LintSeq is an algorithm for computing edit sequence refactorings $\mathcal{D^{\prime}}$ such that all data $(x,\bm{\delta}_{y})\in\mathcal{D^{\prime}}$ have a particular property: every edit in $\bm{\delta}_{y}$ consists of insertions only.
 There are two phases in LintSeq: a backward sampling phase that is used to compute program state sequences $\bm{\sigma}_{y}$, and a forward edit sequence computation phase that is used to re-express $\bm{\sigma}_{y}$ as edit sequences $\bm{\delta}_{y}$.
-An illustration of these phases is shown in Figure 2.
-Full examples of edit sequences generated with LintSeq are provided in Appendix E (Figures 9 and 10).
+Pseudo-code as well as a visualization of each of these phases is provided in Figure 2.
+Full examples of edit sequences generated with LintSeq are provided in Appendix F (Figures 9 and 10).
 Phase I: Backward Sampling
@@
 Starting from the last program that was added to $\bm{\sigma}_{y}$, we use the diff operator to compute edits between each pair of consecutive programs in $\bm{\sigma}_{y}$ up to the original program $y$.
-Finally, we pair each edit sequence $\bm{\delta}_{y}$ with its instruction $x$ (if present) to yield an edit sequence re-factoring $\mathcal{D}^{\prime}$ of $\mathcal{D}$ with size $sN$.
+Finally, we pair each edit sequence $\bm{\delta}_{y}$ with its instruction $x$ (if present) to yield an edit sequence refactoring $\mathcal{D}^{\prime}$ of $\mathcal{D}$ with size $sN$.
 2.4 Properties of LintSeq Data
@@
 In other words, the edit sequence $\bm{\delta}_{y}$ resolves to $y$.
-Furthermore, by construction, every prefix subsequence of $\bm{\delta}_{y}$ resolves to a sub-program of $y$ that is error-free, i.e.
-that throws no errors when checked with the linter or the verifier used during generation.
-These two properties, in conjunction with the uniform sampling step used in the first phase of the algorithm, show that LintSeq samples $s$ examples across all possible static error-free sequences of line insertions that can be used to sequentially write a program $y$ from-scratch.
-We show an example of program synthesis dataset statistics before and after LintSeq processing in Figure 3.
-In the worst case, re-expressing a program as an edit sequence increases the length of a training example by a token count that is constant in the number of program lines (Appendix A).
+Furthermore, by construction, every prefix subsequence of $\bm{\delta}_{y}$ resolves to a intermediate program state of $y$ that is linter-error-free (see Section 2.1).
+These two properties, in conjunction with the uniform sampling step used in the first phase of the algorithm, show that LintSeq samples $s$ examples across all possible linter-error-free sequences of line insertions that can be used to sequentially write a program $y$ from-scratch.
+We show an example of program synthesis dataset statistics before and after LintSeq processing in Appendix A (Figure 6).
+In the worst case, re-expressing a program as an edit sequence increases the length of a training example by a token count that is constant in the number of program lines [^1] See Appendix B for more details.
+.
 2.5 Practicalities of Training Language Models on LintSeq Data
 LintSeq can be run on any code data.
-It is agnostic to the contents of a program, and only depends on knowledge of the programming language that a file is written in and the existence of a linter or another kind of verifier for written program files.
+It is agnostic to the contents of a program, and only depends on knowledge of the language that a program is written in, and the existence of a linter for this language.
 We use teacher-forced supervised learning (Williams & Zipser, 1989) to train models on LintSeq data, concatenating edit sequences into a single string by interleaving edits with special tokens, “<|diff|>,” and computing instruction-conditioned losses over the resultant sequences.
-During inference, finetuned models can be prompted to synthesize programs with edit sequences by appending these special tokens to the ends of prompts.
-More details are provided in Appendix A.
+At test-time, fine-tuned models can be prompted to synthesize programs with edit sequences by appending these special tokens to the ends of prompts.
+More details are provided in Appendix B.
 Synthetic data generation with LintSeq is controlled by a single hyperparameter: the number of edit sequences $s$ that are sampled for each example in the source code dataset $\mathcal{D}$.
-Edit sequence sampling can optionally be constrained to avoid repetitions, though this may yield a dataset $\mathcal{D}^{\prime}$ with fewer than $s|\mathcal{D}|$ examples if there are programs in $\mathcal{D}$ with no non-empty and static error-free sub-programs.
+Edit sequence sampling can be constrained to avoid repetitions.
 3 Experiments
-To study LintSeq and the impact of re-parameterizing program synthesis as a sequential edit generation problem, we conduct a multi-pronged set of instruction finetuning experiments.
+To study LintSeq and the impact of re-parameterizing program synthesis as a sequential edit generation problem, we conduct a set of supervised fine-tuning (SFT) experiments.
 These experiments study code synthesis in Python and are designed to answer the following questions:
-  • How does finetuning tiny LMs on re-factorized code edit data generated with LintSeq impact benchmark coverage compared to finetuning on the original code data?
-  • Do performance improvements hold across model scales, families, and tokenizers?
-  • How does ablating linter-guidance from LintSeq to finetune on randomly sampled edit sequences impact code synthesis?
-Similar to previous works (Chen et al., 2021), we evaluate models by computing zero-shot coverage statistics on code synthesis benchmarks with and without repeated sampling.
-We expound upon our motivation for evaluating models in this manner in Appendix B.3.1.
+  • How does fine-tuning tiny code LMs to generate programs edit-by-edit with supervised learning impact performance on benchmarks compared to fine-tuning on standard code data?
+  • Do performance improvements hold for “off-the-shelf” LMs and on harder coding benchmarks? Do they hold across model scales, tokenizers, and families?
+  • How does ablating linter-guidance from LintSeq impact test-time performance?
+Similar to previous works (Chen et al., 2021), we evaluate models by computing “pass@k,” the probability that at least one of “k” generations for a problem passes all of the unit tests.
 3.1 Pretraining Tiny LMs for Code Understanding
 We begin our investigations by pre-training two tiny decoder-only transformers, TinyCodeLM-150M and TinyCodeLM-400M, for Python code understanding on 72 billion tokens of text.
-Pretraining our own language models grants us a data contamination-free test-bed to study code synthesis with edit sequences, rapidly evaluate LintSeq, and broadly re-examine the trade-off between inference-time compute and generation quality in code synthesis for models that can be updated on-device.
+Pretraining our own language models grants us a data contamination-free test-bed to study code synthesis with edit sequences, rapidly evaluate LintSeq, and broadly re-examine the trade-off between test-time compute and generation quality in code synthesis for models that can be updated on-device.
 We rely on open-source data and libraries to pretrain our models (Penedo et al., 2024; Lozhkov et al., 2024; Soldaini et al., 2024; Groeneveld et al., 2024).
@@
 Our experiments are supported by Pytorch FSDP (Zhao et al., 2023).
-More details on our pretraining procedures are in Appendix C.
+More details on our pretraining procedures are in Appendix D.
 3.2 Generating a Synthetic Dataset with LintSeq
-Next, to support our finetuning experiments, we prepare a large “baseline” dataset of paired instruction and program data.
-We re-factorize the programs in this dataset into code edit sequences with LintSeq.
-Table 1: Summary of temperature-tuned coding benchmark results for LMs with $\mathbf{\leq 0.2}$B parameters.
-Scores annotated with “($\dagger$)” indicate external model evaluations that we ran using the procedure described in Appendix B, and all other scores are as reported by model authors.
-We list models in order of increasing HumanEval pass@1 score.
-HumanEval MBPP(+)
-Model Size pass@1 pass@10 pass@1 pass@10 Open-Source
-PolyCoder 160M 2.1 3.4 - - 🌑
-AlphaCode 89M 4.3 12.2 - - 🌕
-Codex 85M 8.2 12.8 - - 🌕
-SmolLM-Instruct 135M 7.6 14.4($\dagger$) 10.1($\dagger$) 14.6($\dagger$) 🌑
-TinyCodeLM-Instruct 150M 9.1 13.2 11.5 16.4 🌑
-TinyCodeLM-LintSeqInstruct 150M 12.8 20.6 13.6 24.4 🌑
-Table 2: Summary of temperature-tuned coding benchmark results for $\mathbf{0.2}$B to $\mathbf{0.4}$B parameter language models.
-Annotations, model order, and evaluation procedure are the same as in Table 1.
-HumanEval MBPP(+)
-Model Size pass@1 pass@10 pass@1 pass@10 Open-Source
-PolyCoder 400M 3.0 5.3 - - 🌑
-TinyCodeLM-Instruct 400M 9.6 18.5 15.5 22.2 🌑
-SmolLM-Instruct 360M 11.3 19.3($\dagger$) 19.4($\dagger$) 23.1($\dagger$) 🌑
-AlphaCode 302M 11.6 18.8 - - 🌕
-CodeT5+ 220M 12.0 20.7 - - 🌑
-Codegen-Mono 350M 12.8 23.1 9.4($\dagger$) 15.2($\dagger$) 🌑
-Codex 300M 13.2 20.4 - - 🌕
-TinyCodeLM-LintSeqInstruct 400M 13.4 20.9 19.4 29.9 🌑
-To that end, we first pool the Python portions of two open-source instruction datasets for code synthesis: the GPT 3.5/4-based Magicoder instruction tuning dataset and the StarCoder2-15B-based self-alignment training dataset (Wei et al., 2024b; a).
+To support our fine-tuning experiments, we prepare a baseline dataset of paired instruction and program data. We then re-express the programs in this dataset as code edit sequences with LintSeq.
+To that end, we first pool the Python portions of two open-source instruction datasets for code synthesis: the GPT 3.5/4-based Magicoder instruction dataset and the StarCoder2-15B-based self-alignment training dataset (Wei et al., 2024b; a).
 These datasets are generated with the OSS-Instruct approach by Wei et al.
@@
 Furthermore, we strip any chain-of-thought-like natural language explanations from completion data.
-The resultant dataset has over 88,900 instruction-Python program pairs.
-With our baseline dataset prepared, we run LintSeq to generate $s=5$ synthetic edit trajectory samples for each instruction-program pair.
-As described in Section 2.4, we concatenate each synthetic edit trajectory into a single string by interleaving consecutive edits with a special reserved “edit” token.
+The resultant dataset has over 88,900 instruction+program pairs.
+Table 1: HumanEval and MBPP(+) results for TinyCodeLMs after SFT vs existing code models of similar scale ($\leq$ 0.4B parameters).
+Scores annotated with $\dagger$” indicate external model evaluations that we ran using the procedure described in Appendix C, and all other scores are as reported by model authors.
+We list models in order of increasing HumanEval pass@1 and report standard error in computed score.
+Sampling hyperparameters are listed in Appendix C.4.
+HumanEval MBPP(+)
+Model Size pass@1 pass@10 pass@1 pass@10 Open- Source
+AlphaCode 89M 4.3 12.2 - - 🌕
+Codex 85M 8.2 12.8 - - 🌕
+SmolLM-Instruct 135M 7.7 $\pm$ 0.8$\dagger$ 14.5 $\pm$ 1.0$\dagger$ 10.1 $\pm$ 1.8$\dagger$ 14.6 $\pm$ 0.5$\dagger$ 🌑
+TinyCodeLM-Instruct 150M 9.1 $\pm$ 2.3 13.5 $\pm$ 0.6 11.5 $\pm$ 1.9 21.6 $\pm$ 0.4 🌑
+TinyCodeLM-Instruct 400M 11.3 $\pm$ 0.9 18.5 $\pm$ 1.1 15.5 $\pm$ 2.1 22.2 $\pm$ 0.5 🌑
+SmolLM-Instruct 360M 11.3 19.3 $\pm$ 1.1$\dagger$ 19.4 $\pm$ 2.4$\dagger$ 23.1 $\pm$ 0.5$\dagger$ 🌑
+AlphaCode 302M 11.6 18.8 - - 🌕
+CodeT5+ 220M 12.0 20.7 - - 🌑
+TinyCodeLM-LintSeqInstruct 150M 12.8 $\pm$ 2.6 20.6 $\pm$ 1.1 13.6 $\pm$ 2.1 24.4 $\pm$ 0.8 🌑
+Codegen-Mono 350M 12.8 23.1 9.4 $\pm$ 1.8$\dagger$ 15.2 $\pm$ 0.7$\dagger$ 🌑
+Codex 300M 13.2 20.4 - - 🌕
+TinyCodeLM-LintSeqInstruct 400M 13.4 $\pm$ 2.0 20.9 $\pm$ 1.1 19.4 $\pm$ 2.4 29.9 $\pm$ 0.6 🌑
+With our baseline dataset prepared, we run LintSeq to generate $s=5$ synthetic edit sequences for each instruction-program pair.
+As described in Section 2.5, we concatenate each synthetic edit sequence into a single string by interleaving consecutive edits with a special reserved “edit” token.
 Inspired by Muennighoff et al.
@@
 We use the popular Python linter pylint to guide edit sampling during generation.
-Examples of generated edit sequences and experiments testing the effect of varying $s$ are in Appendix E.
-3.3 Finetuning Language Models on LintSeq Edit Sequences
-We now probe the impact of instruction finetuning a variety of autoregressive LMs to synthesize code with edit sequences, compared to standard generation of full programs.
-We use two code synthesis benchmarks to support our model evaluations: HumanEval (Chen et al., 2021) and Mostly Basic Programming Problems (MBPP) (Austin et al., 2021).
-Using both the code edit re-factorized and baseline instruction datasets obtained in section 3.2, we run pairs of finetuning experiments with six different models.
-In each experiment pair, we finetune an LM on both datasets for an equal number of optimizer steps and with the same learning rate schedule, saving intermediate checkpoints throughout finetuning.
-We run full benchmark evaluations on HumanEval and MBPP on each saved checkpoint [^1] To process the generations of edit sequence LMs into executable programs, we simply resolve each of the predicted code edits one-by-one.
-This procedure is visualized in Figure 1 and described in Appendix A.2., performing no prompt tuning.
-Then, we compare the best inference-time scaling behavior, i.e.
-benchmark coverage (pass@k) as a function of samples (k), obtained by edit sequence finetuning vs standard finetuning for each model.
-A more detailed description of the computed metrics as well as a full specification of the evaluation and finetuning procedures is provided in Appendices B and D.
-Figure 4: HumanEval and MBPP coverage with repeated sampling (“pass@k vs.
-k”) achieved by instruction finetuning Gemma 2, Phi-3, and Llama 3.1 language models on a dataset of LintSeq edit sequence re-factored vs standard Python code (temperature $=1$, top-p $=0.95$).
+Examples of generated edit sequences and experiments testing the effect of varying $s$ are in Appendix F.
+3.3 Training Language Models on LintSeq Edit Sequences with SFT
+Next, we probe the impact of training autoregressive LMs to synthesize full programs vs.
+program edit sequences according to natural language instructions.
+Aside from the tiny code LMs described above in Section 3.3.1, we also finetune small LMs from three different model families, ranging in scale from 2.6B to 14B parameters.
+We evaluate tiny code LMs on HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), and small LMs on the additional challenging benchmarks DS-1000 (Lai et al., 2023), BigCodeBench (Zhuo et al., 2024), and CodeContests (Li et al., 2022b).
+Using both the refactored and baseline instruction datasets described in section 3.2, we run pairs of SFT experiments with six different models.
+In each experiment pair, we finetune an LM on both datasets for an equal number of optimizer steps and with the same learning rate schedule, saving intermediate checkpoints throughout fine-tuning.
+Then, we compare the benchmark performance of checkpoints across sampling temperatures [^2] To process the generations of edit sequence LMs into executable programs, we simply resolve each of the predicted code edits one-by-one.
+This procedure is visualized in Figure 1 and described in Appendix B.2., performing no prompt tuning.
+A more detailed description of the computed metrics as well as a full specification of the evaluation and fine-tuning procedures is provided in Appendices C and E.
 3.3.1 TinyCodeLM
-We run our first two pairs of finetuning experiments on TinyCodeLM-150M and TinyCodeLM-400M.
-These models were not pretrained on code synthesis instruction data, nor were they pretrained on any “diff”-like edit data.
-Our experimental results are summarized in Tables 1 and 2, where we compare the temperature-tuned performance of our models to the reported benchmark coverage of existing code LMs at similar parameter scales.
-We also report the inference-time scaling of benchmark coverage as a function of samples for our finetuned models in Appendix Tables 8 and 9.
-For both the 150M and 400M parameter versions of TinyCodeLM, we find that finetuning LMs to synthesize code with edits via LintSeq data dramatically improves benchmark performance compared to baseline finetuning.
-Indeed, the edit sequence variants of TinyCodeLM outperform all existing code language models of comparable scale that we are aware of, including AlphaCode (Li et al., 2022b), Codex (Chen et al., 2021), CodeT5+ (Wang et al., 2023b), and the recent SmolLM-Instruct (Ben Allal et al., 2024).
-Our smaller edit sequence-finetuned model is particularly strong for its size, roughly matching or out-performing models with twice as many parameters including the 300M parameter version of Codex and the 302M-parameter version of AlphaCode (Tables 1 and 2).
+We run our first two pairs of fine-tuning experiments on TinyCodeLM-150M and TinyCodeLM-400M.
+Our experimental results are summarized in Table 1, where we compare the temperature-tuned performance of our models on HumanEval and MBPP(+) to the pass@1 and pass@10 scores of existing LMs with similar parameter counts.
+For both the 150M and 400M parameter versions of TinyCodeLM, we find that fine-tuning LMs to synthesize code with edits via LintSeq data results in stronger benchmark performance compared to the baseline, improving HumanEval pass@1 by 41% ($9.1\mapsto 12.8$) and 19% ($11.3\mapsto 13.4$) and MBPP pass@1 by 18% ($11.5\mapsto 13.6$) and 25% ($15.5\mapsto 19.4$).
+We see a similar scale of improvement on pass@10 for both benchmarks.
+Our smaller LintSeq model is particularly strong for its size, roughly matching the performance of several models with larger parameter counts (Table 1).
+Figure 3: HumanEval, MBPP(+), DS-1000, and BigCodeBench (Instruct) results for Gemma 2, Phi-3, and Llama 3.1 models after SFT on LintSeq (indigo) vs standard Python code (grey).
+On HumanEval and MBPP(+), we tune sampling temp., top-p, and min-p over $\{1,1.1,1.2\}$, $\{0.95,1.0\}$, and $\{0,0.05\}$, respectively with $n=64$ samples.
+On DS-1000, we evaluate models with the completion format, temperature $=0.2$, top-p $=0.5$, min-p $=0$, and $n=40$, following Wei et al.
+(2024b) and Luo et al.
+(2023).
+On BigCodeBench Instruct, we evaluate with greedy decoding (Zhuo et al., 2024).
+Error bars on HumanEval and MBPP scores show standard error.
 3.3.2 Gemma 2, Phi-3, and Llama 3.1
-The results above raise a natural question: do performance improvements from finetuning LMs to synthesize code with edit sequences hold for other model scales, architectures, and tokenizers?
-To test this, we conduct four additional pairs of instruction finetuning experiments on LMs from three model families, Gemma 2, Phi-3, and Llama 3.1, ranging in size from 2.6B to 14B.
-We employ pretrained-only model weights, if available.
-The results of these experiments are in Figure 4, where we plot zero-shot benchmark coverage as a function of samples for instruction finetuned models.
-Raw coverage scores are reported in Appendix Tables 10 and 11.
-Most of our findings echo those of Section 3.3.1.
-Aggregating across sample counts, we find that finetuning models to synthesize code with edits improves overall zero-shot performance on HumanEval and MBPP compared to finetuning on the original data.
-This suggests that re-factoring code with edit sequences is an architecture- and tokenizer-independent mechanism for improving downstream LM outputs.
-Furthermore, as shown in Figure 4 and Tables 10 and 11, we find that the degree by which edit sequence LMs outperform baseline model variants increases with repeated sampling for all tested models, culminating in an average absolute gain in pass@50 of +20% ($\pm$ 3%) on HumanEval and +12% ($\pm$ 2%) on MBPP.
-This observation confirms the hypothesis posed in Section 2, showing that training LMs to synthesize code with edits using LintSeq data improves the relationship between cumulative inference-time compute and zero-shot performance.
-At pass@1, however, our results are slightly more mixed than in Section 3.3.1.
-For Phi-3 models, we observe either no difference or a decrease in score between each pair of model-data variants.
-One explanation for this is bias: the Phi-3 models have been previously instruction finetuned and were likely to have been trained on standard code data, putting LM variants finetuned to generate edit sequence re-factorized code at a comparative disadvantage due to distributional shift.
-Figure 5: HumanEval and MBPP “pass@k vs. k” achieved by finetuning TinyCodeLM models on linter-guided vs randomly sampled code edit sequences (temperature $=1$, top-p $=0.95$).
-3.4 Ablating Linter-Guidance
-The backward sampling phase of LintSeq uses a linter or other verifier to group interdependent lines of a program together during edit sequence generation, ensuring that all code edits resolve to programs that are free of static errors.
-We conclude our experiments by testing the importance of this design choice with TinyCodeLM models.
-To do this, we replace the backwards procedure described in Section 2.2 with exclusively random sampling; during each step of the algorithm, we first sample the number of lines to delete from the current program uniformly at random, before sampling a set of lines with the desired count.
-Using this linter-ablated version of the algorithm, we generate a new synthetic edit sequence dataset with the same size as the LintSeq dataset used in all previous finetuning experiments, i.e.
-with $s=5$ example sequences per sample in the source dataset.
-The average number of edits per example in this dataset ($\overline{E}_{\text{RandSeqInstruct}}=3.9$) is empirically similar to its linter-guided counterpart ($\overline{E}_{\text{LintSeqInstruct}}=3.8$, see Figure 3).
-We employ the same procedure as the one used in Section 3.3 to instruction finetune TinyCodeLM models on the dataset of randomly sampled edit sequences.
-In Figure 5, we compare the temperature $1$ inference-time scaling laws on HumanEval and MBPP obtained by finetuning models on randomly sampled vs static error-free edit sequences.
-Raw model scores are also provided in Appendix F, Tables 8 and 9.
-Ablating linter-guidance results in a decline in benchmark coverage.
-On HumanEval, linter ablation reduces absolute pass@50 score on TinyCodeLM-150M ($22.6\%\mapsto 17.7\%$) and on TinyCodeLM-400M ($26.8\%\mapsto 22.0\%$).
-MBPP pass@50 is similarly affected, dropping for both models ($34.5\%\to 30.2\%$ and $39.6\%\mapsto 34.5\%$).
-These results suggest that the error-free nature of edits in LintSeq instruction finetuning data does indeed have a positive impact on the coding problem coverage of sampled solutions.
-To conclude our analysis, we probe whether training models on error-free edit sequences also has an effect on the presence of errors in across all generated programs, aside from its positive effects on coverage.
-To assess this, we run the Python linter pylint over each of the synthesized programs used to compute the reported temperature $1$ pass@k metrics, checking code for static errors.
-In Figure 6, we plot the total proportions of synthesized program samples with at least one static error across finetuned model variants.
-On both benchmarks, LMs trained on randomly sampled edits (dark grey) appear to generate “buggy” code with much higher frequency than all other models.
-Furthermore, on HumanEval, we find that LintSeq models (indigo) synthesize programs with static errors at a higher frequency than baseline models (light grey), despite their higher coverage of benchmark coding problems.
-This additional finding suggests that model performance gains from LintSeq cannot simply be attributed to improvement in static error frequency across code – training on re-factored code must be helping models write generally better, more diverse programs.
-In summary, the error-free nature of the linter-guided edits sampled in LintSeq appears to indeed be important for improving both the quality and diversity of sampled programs (Figure 5), as well as the overall correctness (Figure 6) of code synthesized by language models trained on edit sequences.
-Figure 6: Comparing static error frequency in synthesized code samples across baseline vs edit sequence instruction finetuned model variants (n $=50$, temperature $=1$, top-p $=0.95$).
+The results above raise a few questions: Do performance improvements from fine-tuning LMs to synthesize code with edit sequences also hold for language models that were not specifically pretrained for code understanding?
+Do they hold across model scales, architectures, and tokenizers?
+To answer these questions, we conduct four additional pairs of SFT experiments on LMs from three model families, Gemma 2, Phi-3, and Llama 3.1.
+We use pretrained-only model weights, if available.
+The selected LMs range in size from 2.6B to 14B and were trained on general-purpose data mixtures (Gemma Team et al., 2024; Abdin et al., 2024; Dubey et al., 2024).
+Our findings align with those presented in Section 3.3.1.
+As shown in Figure 3, LintSeq improves performance on each LMs for all but two of the metrics visualized here (HumanEval pass@1 and BigCodeBench Instruct greedy pass@1).
+Notably, even on these metric, the least performant LintSeq instruction-tuned models still achieve performance that is comparable to the baseline, i.e.
+within standard error of sampling or within a percentage point.
+In aggregate across models, LintSeq improves HumanEval, MBPP, DS-1000, and BigCodeBench Instruct pass@1 by an average absolute gain of +2.3, +4.3, +3.1, and +1.1 in score compared to baseline SFT.
+Furthermore, as shown in Figure 1(right) and Figure 4, the degree by which edit sequence LMs outperform baselines on HumanEval, MBPP, and CodeContests increases with repeated sampling for all tested models.
+In each of the plots included in these figures, we show the total proportion of benchmark problems solved by SFT-ed LMs on any attempt given “k” tries as a function of total test-time compute used during repeated sampling.
+By comparing total test-time compute across model variants, we account for the slight difference between LintSeqInstruct vs Instruct model generation lengths due to the extra “diff” descriptor tokens used by edit sequence models.
+Even after adjusting for these extra tokens, LintSeq consistently improves the relationship between total test-time compute and performance on code synthesis, supporting the hypothesis posed in Section 2.
+In summary, the results of these experiments suggest that refactoring code tuning data into synthetic edit sequences with LintSeq is a code-pretraining-, scale-, architecture-, and tokenizer-independent mechanism for improving the quality and diversity of LM outputs on code generation tasks.
+3.4 Ablating the Linter from LintSeq
+The backward sampling phase of LintSeq uses a linter to decompose code across edits whose contents reflect the syntactical structure of its programming language.
+We conclude our experiments by testing the importance of this design choice with TinyCodeLM models: does fine-tuning on sequences of (entirely) randomly sampled code edits hurt model performance on HumanEval and MBPP(+)?
+Figure 4: Repeatedly sampling from models SFT-ed to generate edit seqs.
+vs full programs: we compare the best pass@k score achieved by modulating sampling hyperparameters for LintSeqInstruct vs Instruct models.
+On HumanEval and MBPP(+), we use the same values as in Figure 3, while on CodeContests, we sweep over temperatures $\{0.5,0.6\}$ and use top-p $=1.0$, min-p $=0$, and $n=128$.
+We then plot benchmark score as a function of the total cost of repeated sampling from each model in FLOPs (see Appendix A.4).
+Shading shows standard error in linear fit.
+See Figure 1 for Phi-3 3.8B and Llama 3.1 8B test-time scaling with repeated sampling curves on HumanEval and MBPP.
+To test this, we replace the backwards procedure described in Section 2.3 with fully random sampling; during each step of the algorithm, we first sample the number of lines to delete from the current program uniformly at random, before sampling a set of lines with the desired count.
+We refer to this algorithm as “RandSeq.” Using RandSeq, we generate a new synthetic edit sequence dataset with the same size as the LintSeq dataset used in all previous fine-tuning experiments.
+The average number of edits per example in this dataset ($\approx 3.9$) is similar to its linter-guided counterpart ($\approx 3.8$) [^3] Note that both datasets also have a similar size in total training tokens ($\approx 18\cdot 10^{6}$ TinyCodeLM tokens)..
+We employ the same procedure as the one used in Section 3.3 to SFT TinyCodeLM models on the RandSeq dataset.
+In Figure 5(left), we compare the pass@1 HumanEval and MBPP score of LintSeqInstruct vs RandSeqInstruct models at high temperatures.
+On both benchmarks and models, ablating the linter from LintSeq hurts performance with statistical significance, reducing HumanEval pass@1 by 30% ($6.4\mapsto 4.5$) and 29% ($8.4\mapsto 6.0$) and MBPP pass@1 by 24% ($8.6\mapsto 6.5$) and 28% ($14.2\mapsto 10.2$), respectively.
+These results suggest that the linter-informed structure of edits in LintSeq fine-tuning data does improve model performance.
+In Figure 5(right), we conclude our analysis by probing whether training models on linted edits has an effect on the total proportion of syntactical errors in completed programs.
+To assess this, we run the Python linter pylint over the full set of generations sampled at temperature $=1$, top-p $=1$, and min-p $=0$, checking each generated program for syntax errors with this linter.
+LMs trained on randomly sampled edits appear to generate “buggy” code with much higher frequency than all other models on both HumanEval and MBPP(+).
+Furthermore, on HumanEval, we find that LintSeq models synthesize programs with linter-errors at a higher frequency than baselines, despite their higher pass@1.
+This additional finding suggests that model performance gains from LintSeq cannot simply be attributed to improvement in low-level correctness of generated code – training on refactored code must be helping models write generally better, more diverse programs.
+Figure 5: Left: HumanEval and MBPP(+) pass@1 achieved by fine-tuning TinyCodeLM models on linter-guided (LintSeq) vs randomly sampled (RandSeq) code edit sequences.
+We tune sampling parameters over the same values as in Figures 3 and 4, and report the best scores for each model.
+Right: Comparing total proportions of generations with lint errors.
+Error bars show standard error.
 4 Related Work
@@
 Unlike Self-Instruct-like synthetic data generation methods, our algorithm does not employ an LLM for data generation, and instead generates examples of error-free edit sequences from existing code data by using a simple linter.
-Finetuning on Edits
-Several works have investigated finetuning code LLMs on edit data.
-Notably, Muennighoff et al.
-(2023) instruction tune models on a 4TB dataset of GitHub commits pairing code changes with human instructions.
+Training on Edits
+Many works have studied edit generation with language models.
+Yin et al.
+(2018) cast the edit representation problem as an autoencoding task and show that neural network models can learn to capture the structure and semantics of edits, while Gu et al.
+(2019) introduce a partially autoregressive model for generating insertion and deletion edits that is trained with adversarial imitation learning.
+Guo et al.
+(2021) use reinforcement learning to train LMs to generate code with “holes” that represent high uncertainty tokens, and to edit the contents of these “holes” later on.
+More recently, several works have investigated finetuning off-the-shelf pre-trained language models on large-scale edit data.
+Berabi et al.
+(2021) use a linter to detect errors in code, and finetune a T5 model (Raffel et al., 2020) to correct code by leveraging error messages.
+Muennighoff et al.
+(2023) and Cassano et al.
+(2023) instruction tune models on datasets of GitHub commits pairing code changes with human instructions.
 Relatedly, Li et al.
 (2024) use GitHub commit data sourced from Python repositories to generate code editing instruction data with GPT 3.5/ChatGPT.
-Both of these works specifically focus on better-equipping LLMs for natural language-prompted code editing tasks, in which a model is explicitly prompted to generate an edit in response to a natural language specification.
-Our work differs in three important ways: first, we study edit sequences rather than single edits; second, we train LLMs to predict edits implicitly during code synthesis; third, our synthetic edit generation algorithm does not rely on the existence of any kind of commit data.
+All of these works specifically focus on better-equipping LMs for natural language-prompted code editing tasks, in which a model is explicitly prompted to generate an edit in response to an error message or a natural language specification.
+Our work differs in three important ways: first, we study edit sequences rather than single edits; second, we train LMs to predict edits implicitly during code synthesis; third, our synthetic edit generation algorithm does not rely on the existence of any kind of commit data.
 “On Device” Language Models
@@
 These models provide an efficient test-bed for experiments on LM code synthesis that is updated to recent advancements in high throughput pretraining and to improvements in open-source data quality.
-Inference-Time Compute Scaling
+Scaling Up Test-Time Compute
 The performance of language models can be boosted during inference by using scaled-up sample counts, hand-engineered prompting schema, and/or search (Brown et al., 2024; Snell et al., 2024).
@@
 This paper introduces an algorithm, LintSeq, for generating synthetic code edit sequences from existing programs.
-LintSeq enables LLM reasoning settings like code synthesis to be re-parameterized at the data-level as sequential edit generation tasks.
-The algorithm is parameter-free, requires only CPU to run, and makes no assumptions about the content or structure of code files.
+LintSeq enables code synthesis to be re-parameterized at the data-level as sequential edit generation tasks.
+The algorithm is parameter-free, requires only CPU to run, and makes no assumptions about the content or structure of source code files.
 Re-parameterizing code generation with edits has a few immediate benefits.
-For example, it makes code generation with LLMs much more controllable at the prompt-level (Appendix A.3) and it reduces the cost of predicting useful and correct code insertions with models, since synthetic edit-trained LLMs do not need to be prompted to re-synthesize entire programs from scratch (Section 2.4).
+For example, it makes code generation with LMs more controllable at the prompt-level (Appendix B.3) and it reduces the cost of predicting useful and syntactically correct code insertions with models, since synthetic edit-trained LMs do not need to be prompted to re-generate full programs from scratch (Section 2.5).
 In our experiments with LintSeq, we also show the following:
   1. 1.
-    Tiny LMs can be efficiently finetuned to synthesize Python programs with edit sequences via LintSeq data.
-This results in state-of-the-art code benchmark performance for models that can be trained “on device” (Sections 3.1 and 3.3.1).
+    Tiny LMs pre-trained for code understanding can be efficiently fine-tuned to synthesize programs edit-by-edit via LintSeq data.
+This results in competitive performance on HumanEval and MBPP(+) compared to existing code LMs of similar scale (Sections 3.1 and 3.3.1).
   2. 2.
-    Across other tested models from the Phi, Gemma, and Llama families, finetuning on LintSeq data also improves the diversity of zero-shot generations, boosting the inference-time scaling of coverage on HumanEval and MBPP at fixed sample counts (Section 3.3.2).
+    On larger models from the Phi 3, Gemma 2, and Llama 3.1 families that were pretrained for general natural language understanding, tuning on LintSeq data either improves or preserves the quality of pass@1 generations compared to standard tuning (Section 3.3.2).
   3. 3.
-    On HumanEval, the cumulative inference cost of repeatedly sampling from small edit sequence LLMs is similar to sampling once from larger LLMs and yields coverage that is competitive with GPT-4, GPT-4-Omni, and Llama 3.1 405B (Figure 1, Appendix F.4).
+    LintSeq also improves test-time compute scaling laws for code synthesis on instruction fine-tuned Phi 3, Gemma 2, and Llama 3.1 models, suggesting that edit sequence LMs consistently generate more meaningfully diverse programs compared to baselines, even on challenging benchmarks like CodeContests (Section 3.3.2).
   4. 4.
-    Ablating linter-guidance from LintSeq hurts the quality, diversity, and correctness of code synthesized by instruction finetuned models (Section 3.4).
+    Ablating the linter from LintSeq hurts the quality and syntactical correctness of code synthesized by edit sequence TinyCodeLMs.
+This suggests that the structured nature of edits sampled with LintSeq is important for downstream LM performance (Section 3.4).
 There are several limitations to our work.
@@
 This is a consequence of the parameter-free nature of the algorithm – every edit in a LintSeq sequence reflects an existing line of code in the source file used to generate it.
-As a result, LintSeq edit data can only be used to train models for synthesis-like tasks, rather than for refinement.
-Second, our experiments with LintSeq study code synthesis in Python only.
-We look forward to testing LintSeq with other programming languages, verifiers, and problems in future work.
-One other exciting setting where LintSeq might improve the tradeoff between model generation quality and inference-time compute is in mathematical reasoning and formal theorem proving.
+As a result, models that are fine-tuned exclusively on data sampled with LintSeq cannot be used for code editing tasks involving deletion edits.
+One simple way to circumvent this limitation might be by mixing LintSeq synthetic edit sequences with human edit data during instruction fine-tuning via datasets like CommitPackFT (Muennighoff et al., 2023), which contain examples of deletions.
+An alternate approach might be to follow-up supervised instruction fine-tuning on LintSeq synthetic data with reinforcement learning in order to train models to interleave insertions with deletions when necessary.
+Second, the experiments that we conducted with LintSeq in this paper studied code synthesis in Python only.
+LintSeq can be similarly used for generating synthetic edit sequences for code written in other programming languages by swapping out the linter using during edit sampling.
+Finally, we used LintSeq to refactor an instruction fine-tuning dataset in this work.
+However, by design, the algorithm can be run on any corpus of source code data, such as The Stack (Kocetkov et al., 2022) or The Stack-v2 (Li et al., 2023).
+In future work, we hope to explore using LintSeq to train LMs to write code edit-by-edit on larger, pre-training scale datasets.
 Ethics Statement
@@
 Any and all model generated code has the potential to be harmful and must not be executed without precautions.
+Reproducibility Statement
+In the supplementary materials accompanying this submission, we provide a Python implementation of LintSeq as well as instructions and code supporting data generation, processing, pretraining, and fine-tuning experiments.
+We also provide thorough textual descriptions of all experimental procedures in the Appendix.
+Appendix C describes prompting and model evaluation, while Appendices D and E detail all of the hyperparameters, procedures, and open-source datasets that we employ for obtaining the results reported throughout Section 3.
+Finally, Appendix A.4 provides references and data for reproducing the results plotted in Figure 1.
 Acknowledgements
@@
 We would like to thank Nate Rahn, Mahi Shafiullah, and David Brandfonbrener for helpful comments and discussions.
-Appendix A More on Edit Sequences and Diffs
-A.1 Reading Unix Diffs
-We provide a guide to reading Unix-style diffs below in Figure 7.
-The diff shown in this figure is computed using the Python library difflib, which is the implementation that we use to compactly represent edits in our synthetic data generation experiments.
-Note that the total extra tokens present in an insertion edit sequence representation of a program scales with the number of program lines $L$, and can be upper-bounded as $T_{\text{diff}}\leq L\cdot((\text{chars in ``decorator''})+(\text{extra chars per line in ``body''}))=16L$.
+Appendix A Additional Results
+A.1 Empirics of Processing Code Data with LintSeq
+Figure 6: Empirics of processing code data with LintSeq.
+Left: Lines per example in a dataset of instruction fine-tuning data for Python synthesis before and after processing with LintSeq via the linter pylint (see Section 3.2).
+LintSeq processing adds lines of diff metadata to examples (see Appendix B).
+Right: The corresponding edit counts per synthetic code edit sequence.
+On a dataset of short programs (14 lines of code, on average), the mean LintSeq edit sequence contains four edits.
+A.2 Comparing LintSeqInstruct to RandSeqInstruct TinyCodeLMs on HumanEval and MBPP(+)
+Table 2: Edit sequence TinyCodeLM results on HumanEval at high sampling temperatures: We tune sampling parameters for edit sequence variants of TinyCodeLM over temperatures (1, 1.1, 1.2), top-p (0.95, 1.0), and min-p (0, 0.05) with $n=64$ completions per problem and report the best pass@k value obtained from each model variant.
+We also report standard error for each score.
+HumanEval
+Model Variant Size Linter Guided pass@1 pass@5 pass@10 pass@20 pass@50
+tinycodeLM-RandSeqInstruct 150M 4.5 $\pm$ 0.4 10.3 $\pm$ 0.5 12.2 $\pm$ 0.5 14.4 $\pm$ 0.6 18.8 $\pm$ 0.6
+tinycodeLM-LintSeqInstruct 150M 6.4 $\pm$ 0.5 13.9 $\pm$ 0.5 16.8 $\pm$ 0.6 19.5 $\pm$ 0.6 23.6 $\pm$ 0.6
+tinycodeLM-RandSeqInstruct 400M 6.0 $\pm$ 0.4 11.7 $\pm$ 0.5 13.9 $\pm$ 0.6 16.4 $\pm$ 0.6 20.8 $\pm$ 0.6
+tinycodeLM-LintSeqInstruct 400M 8.4 $\pm$ 0.4 16.6 $\pm$ 0.6 19.7 $\pm$ 0.6 22.8 $\pm$ 0.6 27.2 $\pm$ 0.6
+Table 3: Edit sequence TinyCodeLM results on MBPP(+) at high sampling temperatures: As above, we tune sampling parameters for all fine-tuned TinyCodeLM variants over temperatures (1, 1.1, 1.2), top-p (0.95, 1.0), and min-p (0, 0.05) with $n=64$ completions per problem and report the best pass@k value obtained from each model variant.
+Standard error is indicated with “$\pm$.”
+MBPP(+)
+Model Variant Size Linter Guided pass@1 pass@5 pass@10 pass@20 pass@50
+tinycodeLM-RandSeqInstruct 150M 6.5 $\pm$ 0.3 17.2 $\pm$ 0.4 22.6 $\pm$ 0.4 27.9 $\pm$ 0.5 34.4 $\pm$ 0.5
+tinycodeLM-LintSeqInstruct 150M 8.6 $\pm$ 0.3 19.5 $\pm$ 0.4 24.5 $\pm$ 0.5 29.0 $\pm$ 0.5 35.1 $\pm$ 0.5
+tinycodeLM-RandSeqInstruct 400M 10.2 $\pm$ 0.4 20.8 $\pm$ 0.4 25.4 $\pm$ 0.5 29.9 $\pm$ 0.5 36.2 $\pm$ 0.5
+tinycodeLM-LintSeqInstruct 400M 14.7 $\pm$ 0.4 25.8 $\pm$ 0.5 29.6 $\pm$ 0.5 33.9 $\pm$ 0.5 39.7 $\pm$ 0.5
+A.3 HumanEval, MBPP(+), CodeContests, DS-1000, and BigCodeBench results for LintSeq vs Baseline Instruction Tuned Gemma 2, Phi-3, and Llama 3.1 Models
+Table 4: Gemma 2, Phi-3, and Llama 3.1 results on HumanEval at high sampling temperatures.
+We report the best pass@k value obtained from each model variant at high sampling temperatures, sweeping over temperature values (1, 1.1, 1.2), top-p (0.95, 1.0), and min-p (0, 0.05).
+We generate $n=64$ completions per problem and report standard error for each estimated score.
+HumanEval
+Model Variant Size pass@1 pass@5 pass@10 pass@20 pass@50
+Gemma-2-Instruct 2.6B 15.3 $\pm$ 0.6 22.0 $\pm$ 0.6 25.2 $\pm$ 0.6 31.6 $\pm$ 0.6 41.7 $\pm$ 0.7
+Gemma-2-LintSeqInstruct 2.6B 22.0 $\pm$ 0.6 34.8 $\pm$ 0.6 41.4 $\pm$ 0.6 48.2 $\pm$ 0.7 55.5 $\pm$ 0.7
+Phi-3-Mini-Instruct 3.8B 35.2 $\pm$ 0.6 49.7 $\pm$ 0.6 55.1 $\pm$ 0.7 59.2 $\pm$ 0.7 62.2 $\pm$ 0.7
+Phi-3-Mini-LintSeqInstruct 3.8B 38.4 $\pm$ 0.6 63.3 $\pm$ 0.6 72.4 $\pm$ 0.6 79.9 $\pm$ 0.6 87.3 $\pm$ 0.5
+Llama-3.1-Instruct 8B 38.4 $\pm$ 0.6 51.3 $\pm$ 0.7 56.2 $\pm$ 0.7 60.2 $\pm$ 0.7 64.2 $\pm$ 0.7
+Llama-3.1-LintSeqInstruct 8B 38.5 $\pm$ 0.6 62.2 $\pm$ 1.6 72.6 $\pm$ 1.6 75.7 $\pm$ 0.6 82.7 $\pm$ 0.6
+Phi-3-Med-Instruct 14B 50.2 $\pm$ 0.6 68.4 $\pm$ 0.6 73.5 $\pm$ 0.6 77.3 $\pm$ 0.6 81.4 $\pm$ 0.6
+Phi-3-Med-LintSeqInstruct 14B 49.7 $\pm$ 0.6 75.0 $\pm$ 0.6 81.6 $\pm$ 0.6 85.9 $\pm$ 0.6 89.6 $\pm$ 0.5
+Table 5: Gemma 2, Phi-3, and Llama 3.1 results on MBPP(+) at high sampling temperatures.
+Exactly as above, we sweep over temperature (1, 1.1, 1.2), top-p (0.95, 1.0), and min-p (0, 0.05) and report the best pass@k value obtained from each model variant.
+We generate $n=64$ completions per problem and report standard error for each estimated score.
+MBPP(+)
+Model Variant Size pass@1 pass@5 pass@10 pass@20 pass@50
+Gemma-2-Instruct 2.6B 20.5 $\pm$ 0.4 30.8 $\pm$ 0.5 34.3 $\pm$ 0.5 37.6 $\pm$ 0.5 41.6 $\pm$ 0.5
+Gemma-2-LintSeqInstruct 2.6B 28.2 $\pm$ 0.5 40.1 $\pm$ 0.5 44.5 $\pm$ 0.5 48.6 $\pm$ 0.5 52.8 $\pm$ 0.5
+Phi-3-Mini-Instruct 3.8B 31.9 $\pm$ 0.5 42.5 $\pm$ 0.5 46.3 $\pm$ 0.5 49.8 $\pm$ 0.5 53.6 $\pm$ 0.5
+Phi-3-Mini-LintSeqInstruct 3.8B 37.2 $\pm$ 0.5 51.4 $\pm$ 0.5 56.1 $\pm$ 0.5 60.3 $\pm$ 0.5 66.0 $\pm$ 0.5
+Llama-3.1-Instruct 8B 37.4 $\pm$ 0.5 50.2 $\pm$ 0.5 53.6 $\pm$ 0.5 56.6 $\pm$ 0.5 60.0 $\pm$ 0.5
+Llama-3.1-LintSeqInstruct 8B 40.3 $\pm$ 0.5 56.2 $\pm$ 0.5 61.1 $\pm$ 0.5 65.5 $\pm$ 0.5 69.4 $\pm$ 0.5
+Phi-3-Med-Instruct 14B 37.7 $\pm$ 0.5 50.4 $\pm$ 0.5 54.0 $\pm$ 0.5 57.0 $\pm$ 0.5 60.1 $\pm$ 0.5
+Phi-3-Med-LintSeqInstruct 14B 39.1 $\pm$ 0.5 55.2 $\pm$ 0.5 60.7 $\pm$ 0.5 65.4 $\pm$ 0.5 71.1 $\pm$ 0.5
+Table 6: Gemma 2, Phi-3, and Llama 3.1 results on CodeContests.
+We sweep over temperature (0.5, 0.6) and use top-p $=1$, min-p $=0$, and $n=128$, and report the best pass@k value obtained from each model variant in the table below.
+We also report standard error for each estimated score.
+CodeContests
+Model Variant Size pass@1 pass@50 pass@100
+Gemma-2-Instruct 2.6B 0.05 $\pm$ 0.05 1.56 $\pm$ 0.26 2.26 $\pm$ 0.30
+Gemma-2-LintSeqInstruct 2.6B 0.61 $\pm$ 0.16 5.71 $\pm$ 0.37 7.03 $\pm$ 0.40
+Phi-3-Mini-Instruct 3.8B 1.80 $\pm$ 0.22 14.86 $\pm$ 0.45 18.59 $\pm$ 0.49
+Phi-3-Mini-LintSeqInstruct 3.8B 2.76 $\pm$ 0.26 19.10 $\pm$ 0.48 22.93 $\pm$ 0.51
+Llama-3.1-Instruct 8B 2.68 $\pm$ 0.28 11.21$\pm$ 0.44 12.80 $\pm$ 0.46
+Llama-3.1-LintSeqInstruct 8B 2.92 $\pm$ 0.27 17.86 $\pm$ 0.47 21.82 $\pm$ 0.51
+Phi-3-Med-Instruct 14B 3.22 $\pm$ 0.27 16.50 $\pm$ 0.47 19.45 $\pm$ 0.50
+Phi-3-Med-LintSeqInstruct 14B 3.02 $\pm$ 0.25 19.09 $\pm$ 0.48 23.11 $\pm$ 0.51
+Table 7: Gemma 2, Phi-3, and Llama 3.1 pass@1 results on DS-1000. We use the same sampling hyperparameters as Luo et al. (2023) and Wei et al. (2024b) to evaluate instruction tuned models.
+Model Variant Size DS-1000, pass@1
+Gemma-2-Instruct 2.6B 2.5
+Gemma-2-LintSeqInstruct 2.6B 3.8
+Phi-3-Mini-Instruct 3.8B 8.6
+Phi-3-Mini-LintSeqInstruct 3.8B 15.5
+Llama-3.1-Instruct 8B 14.5
+Llama-3.1-LintSeqInstruct 8B 16.2
+Phi-3-Med-Instruct 14B 21.8
+Phi-3-Med-LintSeqInstruct 14B 24.2
+Table 8: Gemma 2, Phi-3, and Llama 3.1 pass@1 results on BigCodeBench (Instruct). We use greedy decoding to evaluate instruction tuned models.
+Model Variant Size BigCodeBench Instruct, pass@1
+Gemma-2-Instruct 2.6B 5.44
+Gemma-2-LintSeqInstruct 2.6B 6.32
+Phi-3-Mini-Instruct 3.8B 20.79
+Phi-3-Mini-LintSeqInstruct 3.8B 21.58
+Llama-3.1-Instruct 8B 21.46
+Llama-3.1-LintSeqInstruct 8B 20.53
+Phi-3-Med-Instruct 14B 24.65
+Phi-3-Med-LintSeqInstruct 14B 28.16
+A.4 Computing Pass@k vs Total Test-Time FLOPs
+In Figures 1(right) and 4, we plot the percentage of problems solved by any attempt (i.e.
+pass@k) on HumanEval, MBPP, and CodeContests as a function of total test-time FLOPs used during sampling for LintSeq vs baseline instruction fine-tuned models.
+Raw “pass@k” estimates are also included in Tables 4, 5, and 8, representing the best scores achieved by each model variant after tuning sampling hyperparameters.
+We compute total test-time FLOPs using the approximations below, which are drawn from Kaplan et al.
+(2020).
+These approximations conservatively estimate the cumulative inference costs of synthesizing solutions to all of the problems in the test set of each benchmark.
+The models that we compare are all dense transformers, where the majority of the parameters are used in matrix multiplications.
+FLOPs per token $\approx 2\cdot(N_{\text{model-params}}+2\cdot L_{\text{model-layers}}\cdot C_{\text{context}})$
+Total FLOPs $\approx\text{FLOPs per token}\cdot T_{\text{avg-total-tokens-per-sample}}\cdot K_{\text{samples}}\cdot M_{\text{problems}}$
+We determine the quantities $T_{\text{avg-total-tokens-per-sample}}$ for each model variant at a particular “pass@k” by computing token counts over all sets of samples per problem.
+Note that edit sequence (i.e. LintSeqInstruct fine-tuned) LMs have slightly higher average token counts per sample due to presence of “diff” descriptor tokens in generations (see Appendix B).
+Appendix B More on Edit Sequences and Diffs
+B.1 Reading Unix Diffs
+[…]
 Figure 7: The anatomy of a Unix diff: A diagrammatic visualization of the different parts of a Unix-style diff, as computed by difflib.
 The body of a diff can consist of multiple line deletions, followed by multiple line insertions.
-The decorator portion of the diff respectively indicates the location and size of these deletions and insertions, if any.
+The decorator portion of the diff shows the location and size of these deletions and insertions, if any.
 Like the diff shown above, the edits in synthetic edit sequences generated by LintSeq consist of line insertions only.
-A.2 Resolving Edit Sequences
-During inference, LMs that have been finetuned on LintSeq instruct data will synthesize code via edit sequences, outputting text strings that consist of a sequence of consecutive Python diffs interleaved with newline characters and “<|diff|>” tokens, similar to  Piterbarg et al.
+B.2 Resolving Edit Sequences
+During inference, LMs that have been fine-tuned on LintSeq instruct data will iteratively synthesize programs by generating edits i.e., outputting text that consists of a sequence of consecutive Python diffs interleaved with newline characters and “<|diff|>” tokens, similar to  Piterbarg et al.
 (2024).
-Each of these diffs will be structured as shown in Figure 7, if correctly formatted by the language model.
+If correctly formatted by the LM, these diffs will be structured as shown in Figure 7.
 Resolving an edit sequence generated by a language model into an executable Python program is simple: starting with an empty program, we consecutively apply the line insertions and/or deletions in the body of each diff to the lines of the program specified in its decorator.
 We continue this process until all of the diffs in the generated edit sequence have been parsed and resolved.
-Figure 1 shows a code edit sequence generation from a LintSeq instruction finetuned LM and the corresponding resolved, executable Python program.
-A.3 Controllability of Code Synthesis with Edit Sequence LMs
+Figure 1 shows a code edit sequence generation from a LintSeq instruction fine-tuned LM and the corresponding resolved, executable Python program.
+B.3 Controllability of Code Synthesis with Edit Sequence LMs
 The structure of Unix-style diffs affects the downstream controllability of code synthesis with models that have been trained on edit sequence re-parameterized programs.
-As shown in Figure 7, the first line of every diff is a decorator that describes the location and the numbers of lines changed by the edit.
-During inference, autoregressive language models that have been trained on Unix-style diffs with this format can be prompted to predict an edit in any desired target location within the program being synthesized by “intervening” on a model generation.
-A.4 Future Work: Searching in Edit Space
-If we apply the lens of reinforcement learning or search to this setting, we might say that re-parameterizing the code data used to train a language model re-parameterizes the model’s generative action space.
-It is possible that combining edit sequence LMs with more sophisticated decoding mechanisms, inference-time search, and/or interactive post-training may result in even larger improvements to the quality of generated code than those of the zero-shot code synthesis settings studied in this paper.
-We look forward to testing this hypothesis in future work.
-Appendix B Evaluation
+As shown in Figure 7, the first line of every diff is a decorator that describes the location and the number of lines changed by the edit.
+During inference, autoregressive language models that have been trained on diffs with this format can be prompted to predict an edit in a target location by intervening on a model generation.
+B.4 Future Work: Searching in Edit Space
+If we apply the lens of reinforcement learning or search to this setting, we might say that re-parameterizing the code data used to train a language model re-parameterizes the model’s action space.
+It is possible that combining edit sequence LMs with more sophisticated decoding mechanisms, test-time search, and/or reinforcement learning may result in even larger improvements to the quality of generated code than those of the zero-shot code synthesis settings studied in this paper.
+We look forward to testing this in future work.
+Appendix C Evaluation
 HumanEval (Chen et al., 2021) and Mostly-Basic Programming Problems (MBPP) (Austin et al., 2021) are two of the most studied benchmarks for evaluating code LMs (Liu et al., 2023).
@@
 (2023) to add additional rigour to our testing procedure.
-All of the code LMs that we compare our models against evaluate HumanEval performance using the original set of benchmark test cases; for consistency, we employ these same test cases in our evaluations when comparing the performance of our models to the reported scores of external LMs.
-During testing on both HumanEval and MBPP(+), LMs are prompted to generate outputs using the natural language descriptions of target programs.
+The code LMs that we compare our TinyCodeLM models against in Table 1 evaluate HumanEval performance using the original set of benchmark test cases; for consistency, we employ these same test cases in all of our evaluations.
+Our evaluations on the harder benchmarks CodeContests, DS-1000, and BigCodeBench(Instruct) use exactly the same sets of problem descriptions and test cases as those introduced by Li et al.
+(2022b), Lai et al.
+(2023), and Zhuo et al.
+(2024).
+During testing on each benchmarks, LMs are prompted to generate outputs using the natural language descriptions of target programs.
 Their outputs are then evaluated on the paired test cases.
 A generation is considered “correct” if and only if it passes all of the test cases upon execution, subject to a fixed timeout setting.
-Previous works on code synthesis with language models report scores on HumanEval and MBPP(+) across samples.
+Previous works on code synthesis with language models report scores across samples.
 The most common of these metrics is known as pass@k (Chen et al., 2021; Austin et al., 2021; Li et al., 2022b; Wang et al., 2023b).
 This is the metric that we use to report and compare model performance throughout this paper.
-Our implementation of HumanEval and MBPP(+) evaluations mimics the Big Code Evaluation Harness by Ben Allal et al.
-(2022).
-We do not allow models to use chain-of-thought during generation.
-B.1 Prompting
-The primary goal of this paper is to introduce a method for re-factorizing code synthesis with LMs by finetuning them on synthetic instruction data.
+C.1 Prompting
+The primary goal of this paper is to introduce a method for re-factorizing code synthesis with LMs by fine-tuning them on synthetic instruction data.
 As a result, we evaluate all models using minimal prompt formats, performing no prompt tuning (see Figures 9 and 10).
@@
 To evaluate models on HumanEval, we use both the default “Python version” prompt format in the original benchmark dataset, where a natural language program description is provided to an LM within a docstring, as well as the equivalent, fully natural language prompt format from HumanEvalPack (Muennighoff et al., 2023).
-The latter format is similar to the structure of the instructions in our finetuning datasets.
+The latter format is similar to the structure of the instructions in our fine-tuning datasets.
+We report results on the prompt format that yields the best score for each model.
 To evaluate models on MBPP(+), we use the default prompts from the MBPP benchmark dataset, formatted with specification of the target function name and arguments both inside and outside of the natural language instruction, as shown in Figure 8.
-During LM benchmark evaluations, we test models on each of the format variants described above and report scores on the better performing variant for each benchmark only.
-B.2 Generation and Parsing
+As on HumanEval, we report results on the prompt format that yields the best score for each model.
+To evaluate models on BigCodeBench(Instruct) and CodeContests, we simply prompt models with the problem descriptions introduced in the original version of the benchmark (Zhuo et al., 2024; Li et al., 2022b).
+Finally, to evaluate models on DS-1000, we use the completion format, with precisely the same prompt structures as those used by Wei et al. (2024b).
+C.2 Generation and Parsing
 During generation, we continue decoding until an end-of-sequence token is output by an LM.
-We treat all LM outputs as either Python code or sequences of Python code edits, depending on whether an LM was finetuned on standard instruct or LintSeq instruct data.
-In the latter case, we post-process outputs by resolving the output edit sequences using the procedure described in Appendix A.2.
-B.3 Evaluating Model Checkpoints
-B.3.1 Philosophy
+We treat all LM outputs as either Python code or sequences of Python code edits, depending on whether an LM was fine-tuned on standard instruct or LintSeq instruct data.
+In the latter case, we post-process outputs by resolving the output edit sequences using the procedure described in Appendix B.2.
+C.3 Evaluating Model Checkpoints
+C.3.1 Philosophy
 There is a well-known trade-off between the temperature used for sampling from autoregressive code LMs and the benchmark coverage achievable by models, i.e.
@@
 […]
-B.3.2 Computing Coverage (Pass@K)
-Our goal is to probe whether re-parameterizing code synthesis with edit sequences can improve the expressivity of smaller LLM representations, boosting benchmark coverage as a function of samples-per-problem.
-Hence, we primarily compare finetuned models by evaluating them with the procedures described above on HumanEval and MBPP(+) at a high temperature and a large sample count, computing pass@k for k$\ \in\{1,5,10,20,50\}$ with $N=50$ samples [^2] These are the largest sample counts that are feasible to compute on our hardware given the scope of our experiments.
-at temperature 1, top-p 0.95.
-We compute pass@k statistics with the same procedure as Chen et al.
+C.3.2 Computing Pass@K
+Our goal is to probe whether re-parameterizing code synthesis with edit sequences can improve the expressivity of smaller LM representations, boosting benchmark scores as a function of total test-time compute.
+Hence, we primarily compare fine-tuned models by evaluating them with the procedures described above across multiple pass@k.
+We compute unbiased pass@k statistics with the same procedure as Chen et al.
 (2021).
-The results of these evaluations are reported throughout the paper and shown in Figures 4, 5 and Tables 8, 9, 10, 11.
-In each of these Figures and Tables, we identify the most performant checkpoint from each model-data finetuning run by comparing pass@50 score at temperature 1 on HumanEval and MBPP(+) across checkpoints.
+The results of these evaluations are reported throughout the paper.
+C.4 Comparing TinyCodeLMs to existing Models in Table 1
 Many existing state-of-the-art code synthesis LMs only report temperature-tuned pass@k scores on HumanEval, including Codex, AlphaCode, and Codegen-Mono (Chen et al., 2021; Li et al., 2022b; Nijkamp et al., 2022).
-Thus, in Tables 1 and 2, we select the single best overall checkpoint of each TinyCodeLM model by averaging HumanEval and MBPP(+) pass@50 score.
-Then, we temperature-tune each checkpoint’s pass@1 and pass@10 scores when reporting results.
+Thus, in Table 1, we temperature-tune TinyCodeLM models’ pass@1 and pass@10 scores when reporting results.
 On HumanEval, we test temperatures $\tau\in\{0.0,0.2,0.4,0.8,1.0\}$.
@@
 We perform the same temperature tuning procedure when reporting external model benchmark scores as well, i.e.
-the scores annotated with “()($\dagger$)( † )” in Tables 1 and 2.
-When running benchmark evaluations with these external code LMs, we stray from the prompt formatting, generation, and parsing procedures described in Appendices B.1 and B.2; instead, in the interest of a fair evaluation, we reproduce the conventions reported by model authors to report other scores.
-Appendix C Pretraining
+the scores annotated with “()($\dagger$)( † )” in Table 1.
+When running benchmark evaluations with these external code LMs, we stray from the prompt formatting, generation, and parsing procedures described in Appendices C.1 and C.2; instead, in the interest of a fair evaluation, we reproduce the conventions reported by model authors to report other scores.
+Appendix D Pretraining
 We rely on data and libraries open-sourced by the HuggingFace, FineWeb, StarCoder, Dolma, OLMo, and PyTorch FSDP projects to pretrain our models (Wolf et al., 2020; Penedo et al., 2024; Lozhkov et al., 2024; Soldaini et al., 2024; Groeneveld et al., 2024; Zhao et al., 2023).
-C.1 Model Architectures and Pretraining Hyperparameters
+D.1 Model Architectures and Pretraining Hyperparameters
 Table 9: Architectural and pretraining hyperparameters of our “on device” 150M and 400M parameter TinyCodeLM models, pretrained on a mixture of Web text and code for Python understanding.
@@
 Total Epochs of Pretraining 2 2
-C.2 Pretraining Data Mix
+D.2 Pretraining Data Mix
 Table 10: Pretraining data mix used to train both TinyCodeLM models. Datasets were tokenized and prepared using HuggingFace and Dolma tooling (Wolf et al., 2020; Soldaini et al., 2024).
@@
 The Stack (Kocetkov et al., 2022) Python Only 61.8BT 24.2M
-Appendix D Instruction Finetuning
-D.1 Baseline Instruction Dataset
+Appendix E Instruction Fine-tuning
+E.1 Baseline Instruction Dataset
 Table 11 displays the data sources that are used to prepare the dataset described in Section 3.2.
 These data are pooled and preprocessed into instruction-program pairs by stripping away Markdown formatting and natural language explanations from completions (Figure 9 and 10).
-In our experiments, we use the resultant data to finetune baseline models, comparing their performance to those of LMs finetuned on edit sequences generated with LintSeq from the same set of instruction-program pairs.
+In our experiments, we use the resultant data to finetune baseline models, comparing their performance to those of LMs fine-tuned on edit sequences generated with LintSeq from the same set of instruction-program pairs.
 HuggingFace Instruction Data Source Subset Examples
@@
 ise-uiuc/Magicoder-OSS-Instruct-75K Python 38,284
-88,945
 Table 11: Instruction data mix used to prepare the baseline instruction dataset in Section 3.2.
-D.2 Procedures and Hyperparameters
+E.2 Procedures and Hyperparameters
 We instruction finetune all models with Microsoft DeepSpeed using the ZeRO++ protocol for stage three sharding.
 For the largest of these models, we also use CPU parameter offloading to accelerate experiments (Wang et al., 2023a; Ren et al., 2021).
-When finetuning models on LintSeq data, we add a new token “<|diff|>” to tokenizers (Section 2.4) and resize model embeddings accordingly.
+When fine-tuning models on LintSeq data, we add a new token “<|diff|>” to tokenizers (Section 2.5) and resize model embeddings accordingly.
 In our experiments with Gemma 2, Phi-3, and Llama 3.1 models, we use HuggingFace to access and load pretrained model weights and tokenizers.
@@
 This is the case for Gemma 2 and Llama 3.1 only, as of the writing of this paper.
-Across all of the finetuning experiments conducted in this paper, we train model-data variants with the same batch size and for an equal number of total optimizer steps.
-This optimizer step count corresponds to ten epochs of finetuning with the baseline instruction tuning dataset described in Section 3.2.
-We save intermediate checkpoints at equal optimizer step intervals in all experiments.
+Across all of the fine-tuning experiments conducted in this paper, we train model-data variants with the same batch size and for an equal number of total optimizer steps.
+This optimizer step count corresponds to ten epochs of fine-tuning with the baseline instruction tuning dataset described in Section 3.2.
+We save intermediate checkpoints at equal optimizer step intervals in all experiments, and we report benchmark scores for the best performing checkpoint from each model-data variant.
 […]
@@
 Learning Rate Scheduler linear
+Max Learning Rate 1e-4
 Warmup Ratio 0.001
@@
 Total Optimizer Steps 1740
-Table 7: All other instruction finetuning settings, re-used across experiments.
-Appendix E More on Synthetic Data Generation with LintSeq
-E.1 Examples of Generated Synthetic Edit Trajectories
+Table 13: All other instruction fine-tuning settings, re-used across experiments.
+Appendix F More on Synthetic Data Generation with LintSeq
+F.1 Examples of Generated Synthetic Edit Trajectories
 Figure 9: LintSeq edit sequence samples vs baseline instruction-program data, example A.
 Figure 10: LintSeq edit sequence samples vs baseline instruction-program data, example B.
-E.2 Tuning LintSeq Example Count
+F.2 Tuning LintSeq Example Count
 Figure 11: Probing the effect of varying the number of edit sequences sampled with LintSeq per instruction-example pair during data generation: Using the source dataset described in Section 3.2, we sweep over the value of the LintSeq parameter $s$ used during synthetic data generation to yield three different edit sequence instruction datasets with $s\in\{1,5,10\}$.
@@
 The most performant values is $s=5$.
-Appendix F Additional Results
-F.1 Pretraining TinyCodeLM
-Figure 12: Evaluating the zero-shot Python synthesis capabilities of TinyCodeLM-150M during pretraining on HumanEval and MBPP(+).
-Figure 13: Evaluating the zero-shot Python synthesis capabilities of TinyCodeLM-400M during pretraining on HumanEval and MBPP(+).
-F.2 Finetuning TinyCodeLM
-F.2.1 Inference-Time Scaling Laws
-Table 8: HumanEval fixed-temperature coverage scaling results achieved by finetuning TinyCodeLM models (zero-shot, temperature = 1, top-p = 0.95).
-HumanEval
-Model Variant Size Linter Guided pass@1 pass@5 pass@10 pass@20 pass@50
-tinycodeLM-Instruct 150M - 6.2 10.4 12.3 14.5 18.3
-tinycodeLM-RandSeqInstruct 150M 4.0 9.7 11.9 14.4 17.7 (-0.6)
-tinycodeLM-LintSeqInstruct 150M 6.4 14.3 17.7 20.4 22.6 (+4.3)
-tinycodeLM-Instruct 400M - 6.8 11.4 13.7 16.2 18.9
-tinycodeLM-RandSeqInstruct 400M 7.2 12.7 15.6 18.6 22.0 (+3.1)
-tinycodeLM-LintSeqInstruct 400M 7.2 14.9 18.2 21.5 26.8 ( +7.9)
-Table 9: MBPP(+) fixed-temperature coverage scaling results achieved by finetuning TinyCodeLM models (zero-shot, temperature = 1, top-p = 0.95).
-MBPP(+)
-Model Variant Size Linter Guided pass@1 pass@5 pass@10 pass@20 pass@50
-tinycodeLM-Instruct 150M - 7.3 16.4 20.4 24.6 30.6
-tinycodeLM-RandSeqInstruct 150M 4.3 13.1 18.4 23.7 30.2 (-0.4)
-tinycodeLM-LintSeqInstruct 150M 7.5 18.2 23.4 28.5 34.5 (+3.9)
-tinycodeLM-Instruct 400M - 9.3 17.9 22.2 26.8 32.4
-tinycodeLM-RandSeqInstruct 400M 5.5 15.9 21.5 27.0 34.5 (+2.1)
-tinycodeLM-LintSeqInstruct 400M 13.1 25.2 29.9 34.4 39.6 (+7.2)
-F.3 Finetuning Gemma 2, Phi-3, and Llama 3.1
-F.3.1 Inference-Time Scaling Laws
-Table 10: HumanEval fixed-temperature coverage scaling results achieved by finetuning Gemma 2, Phi-3, and Llama 3.1 models (zero-shot, temperature = 1, top-p = 0.95).
-HumanEval
-Model Variant Parameters pass@1 pass@5 pass@10 pass@20 pass@50
-Gemma-2-Instruct 2.6B 16.1 26.3 30.5 34.6 40.2
-Gemma-2-LintSeqInstruct 2.6B 21.6 35.0 41.2 48.3 58.5 (+18.3)
-Phi-3-Mini-Instruct 3.8B 40.1 48.8 52.1 55.0 57.3
-Phi-3-Mini-LintSeqInstruct 3.8B 38.3 63.6 72.0 79.0 86.6 (+29.3)
-Llama-3.1-Instruct 8B 34.2 50.3 55.6 59.7 63.4
-Llama-3.1-LintSeqInstruct 8B 37.6 61.1 69.2 75.3 82.3 (+18.9)
-Phi-3-Med-Instruct 14B 52.8 67.9 72.0 75.0 77.4
-Phi-3-Med-LintSeqInstruct 14B 44.7 73.5 80.2 84.5 90.2 (+12.8)
-Table 11: MBPP(+) fixed-temperature coverage scaling results achieved by finetuning Gemma 2, Phi-3, and Llama 3.1 models (zero-shot, temperature = 1, top-p = 0.95).
-MBPP(+)
-Model Variant Parameters pass@1 pass@5 pass@10 pass@20 pass@50
-Gemma-2-Instruct 2.6B 22.8 34.2 37.8 41.2 44.6
-Gemma-2-LintSeqInstruct 2.6B 27.7 41.3 46.3 50.5 54.7 (+10.1)
-Phi-3-Mini-Instruct 3.8B 35.2 45.9 49.1 52.0 56.1
-Phi-3-Mini-LintSeqInstruct 3.8B 39.3 57.5 62.4 65.9 69.4 (+13.3)
-Llama-3.1-Instruct 8B 38.4 50.2 53.8 56.6 59.4
-Llama-3.1-LintSeqInstruct 8B 38.5 56.5 61.6 65.7 69.8 (+10.4)
-Phi-3-Med-Instruct 14B 40.3 51.5 54.2 56.3 59.0
-Phi-3-Med-LintSeqInstruct 14B 40.0 60.9 67.0 71.5 75.9 (+16.9)
-F.4 Computing HumanEval Coverage vs Cumulative Inference-Time FLOPs
-In Figure 1, we plot HumanEval coverage as a function of cumulative inference-time FLOPs, comparing the performance and total cost of repeatedly sampling from our instruction finetuned Phi-3 and Llama 3.1 models vs sampling a single generation per problem from very large models like Llama 3.1 405B (Dubey et al., 2024) and Nemotron 4 340B (Adler et al., 2024).
-We use the approximations below, drawn from Kaplan et al.
-(2020), to conservatively estimate the cumulative inference costs of synthesizing solutions to all of the $164$ HumanEval benchmark problems across different models.
-The models that we compare are all dense transformers, where the majority of the parameters are used in matrix multiplications.
-FLOPs per token $\approx 2\cdot(N_{\text{model-params}}+2\cdot L_{\text{model-layers}}\cdot C_{\text{context}})$
-Total FLOPs $\approx\text{FLOPs per token}\cdot T_{\text{avg-total-tokens-per-sample}}\cdot K_{\text{samples}}\cdot M_{\text{problems}}$
-$\approx\text{FLOPs per token}\cdot T_{\text{avg-total-tokens-per-sample}}\cdot K_{\text{samples}}\cdot 164$
-For our instruction finetuned models, we determine the quantities $T_{\text{avg-total-tokens-per-sample}}$ by computing token counts over all sets of samples per problem that we obtained to compute the coverage statistics in Figure 4 and Table 10 above.
-These token statistics are provided in the table below.
-Table 12: HumanEval total tokens per sample for finetuned Gemma 2, Phi-3, and Llama 3.1 models (zero-shot, temperature $=1$, top-p $=0.95$).
-These counts reflect prompt and completion tokens.
-They are computed using the same samples whose coverage statistics are reported in Table 10.
-Model Variant Parameters Avg. Total Tokens Per HumanEval Sample ($n=50$)
-Gemma-2-Instruct 2.6B 172
-Gemma-2-LintSeqInstruct 2.6B 205
-Phi-3-Mini-Instruct 3.8B 218
-Phi-3-Mini-LintSeqInstruct 3.8B 247
-Llama-3.1-Instruct 8B 178
-Llama-3.1-LintSeqInstruct 8B 201
-Phi-3-Med-Instruct 14B 230
-Phi-3-Med-LintSeqInstruct 14B 247
-Note that edit sequence (i.e. LintSeqInstruct finetuned) LMs have slightly higher average token counts per sample due to presence of “diff” descriptor tokens in generations (see Appendix A).
-We report zero-shot HumanEval coverage for external models by using the evaluation statistics from Dubey et al.
-(2024) (Table 18, column one).
-To estimate cumulative inference-time FLOPs for these models, we employ the approximation expression above and estimate $T_{\text{avg-total-tokens-per-sample}}\approx 200$, reflecting an ensemble over the per sample token counts of standard “Instruct” finetuned models shown in Table 12.
-Note that this ensembled statistic reflects program generations without chain-of-thought only.
-As a result, we believe it to be a conservative estimator.